### v3

# PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection

Abstract
Anomaly detection stands as a crucial aspect of time series analysis, aiming to identify abnormal events in time series samples. The central challenge of this task lies in effectively learning the representations of normal and abnormal patterns in a label-lacking scenario. Previous research mostly relied on reconstruction-based approaches, restricting the representational abilities of the models. In addition, most of the current deep learning-based methods are not lightweight enough, which prompts us to design a more efficient framework for anomaly detection. In this study, we introduce PatchAD, a novel multi-scale patch-based MLP-Mixer architecture that leverages contrastive learning for representational extraction and anomaly detection. Specifically, PatchAD is composed of four distinct MLP Mixers, exclusively utilizing the MLP architecture for high efficiency and lightweight architecture. Additionally, we also innovatively crafted a dual project constraint module to mitigate potential model degradation. Comprehensive experiments demonstrate that PatchAD achieves state-of-the-art results across multiple real-world multivariate time series datasets. Our code is publicly available.  [^1] https://github.com/EmorZz1G/PatchAD
1 Introduction
Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from normal behavior Belay et al. (2023); Choi et al. (2021). This process becomes more complex in multivariate temporal anomaly detection, where it involves multiple channel dimensions. Driven by the rapid progression of large-scale sensing technologies and enhanced data storage capabilities, temporal anomaly detection has found widespread application in various sectors. For instance, in the field of the Internet of Things (IoT), it serves to monitor abnormal events in sensor data Yang et al. (2021). In healthcare sector, it plays a vital role in monitoring a person’s physiological data, thereby assisting in the early detection of health problems Cook et al. (2020).
However, extracting anomalies from extensive and complex temporal data presents significant challenges. First, accurately defining the representations of anomalies is difficult due to the varied and evolving nature of real-world anomalies. For instance, anomalies in machinery could manifest as either high or fluctuating temperatures under different scenarios. Second, labels for anomalous events in time-series data are probably absent or hard to acquire. Thirdly, the high dimensionality and temporal dependencies inherent in time-series data are intricate. While the typical approach involves classifying different time points as normal or anomalous, relying solely on a single time point is often not feasible.
Due to the powerful representational capacity of neural networks, numerous deep learning-based techniques Munir et al. (2019); Zhang et al. (2022) have been proposed for time series anomaly detection, demonstrating remarkable performance and attracting significant interest from the research community. While supervised and semi-supervised methods grapple with the scarcity of labeled data, unsupervised techniques Zhou et al. (2019); Su et al. (2019) have gained prominence due to their ability to operate without strict labeling requirements. While Xu et al. (2022) designed six different anomaly injection methods to train the model in a self-supervised approach, Carmona et al. (2021) utilized outlier exposure and mixup to construct a self-supervised task.
Reconstruction-based approaches, as typical and common unsupervised methods, involve learning representations of historical data and identifying samples that cannot be accurately reconstructed during testing as anomalies Park et al. (2018). This approach has experienced rapid development due to its clear concept and its proficiency in managing complex data representations. However, due to their robust reconstruction capabilities, these methods excel at recovering anomalous data, posing a challenge for reconstruction-based approaches. The core issue is the ineffectiveness of these methods in distinctly modeling normal and abnormal data, complicating their ability to accurately identify anomalies.
Recently, contrastive learning has been captivated by its straightforward, yet exceptional performance, proving effective in areas such as computer vision Chen et al. (2020). In time series anomaly detection, several methods have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures tailored for contrastive learning. They either do not explicitly model inter-channel dependencies Yang et al. (2023) or rely solely on basic MLPs or Transformers that focus on temporal information Zhou et al. (2022), mainly considering point-level features Audibert et al. (2020). Nevertheless, poor training and inference speeds, architectural complexity, and a lack of lightweight features are some of the drawbacks of high-performing techniques. It was recently shown by Zeng et al. (2022) that MLP structures can also result in outstanding performance on time series tasks. Additionally, they neglected the model degradation issue in contrastive learning, meaning models might degrade or collapse during the training process.
In this paper, we propose PatchAD, a novel Patch-based MLP architecture designed for time series Anomaly Detection. PatchAD leverages contrastive learning to tackle the complexities of time series anomaly detection. Inspired by Yang et al. (2023), we hypothesize that normal time series points share similar latent patterns, indicating a strong correlation among normal time points. On the contrary, anomalies struggle to establish connections with other anomaly points or normal points. Hence, normal data exhibits consistent representations across different views, while anomalies struggle to maintain consistent representations. This underlying principle forms the core of our approach to anomaly detection.
Specifically, we introduce a novel Multi-scale Patching and Embedding approach. This method decomposes the input data along the temporal dimension into non-overlapping patches, enabling the model to learn from features that contain richer semantic information, moving beyond simple point-level features. To address various time scales, different submodels are utilized, each focusing on different patch sizes. To capture diverse views of temporal data representation, we propose different Patch Mixer Encoders to learn the mixed relationships among inter-patches, intra-patches, and inter-channel connections. We utilize a shared MixRep Mixer to uniformly represent the feature space. To prevent the degradation, we not only construct diverse encoders but also introduce a Dual Project Constraint mechanism to prevent the model from settling on overly simplistic solutions, thus substantially boosting its representational capabilities.
The main contributions of our proposed PatchAD can be summarized as follows:
  • The novel PatchAD model, innovatively structured on contrastive learning framework, is adept at managing diverse anomaly types. This is achieved by creating Interpatch Encoders and Intra-patch Encoders, which generate both inter-patch and intra-patch perspectives.
  • Employing a strategy named Multi-scale Patching and Embedding, PatchAD integrates multi-scale subsequence modules constructed exclusively with MLPs. This design is intended to facilitate the learning of semantically richer features.
  • Dual Project Constraint mechanism is specifically designed to impose constraints on the model, reducing its susceptibility to simple, trivial solutions and thus increasing its expressive power.
  • PatchAD’s efficacy is underscored by comprehensive experimental results on eight datasets. These results highlight its superior performance compared to the SOTA time series anomaly detection methods.
Figure 1: The workflow of the proposed PatchAD framework.
2 Preliminaries
In this section, we provide an initial exposition on the formal representation of time series anomaly detection, related methods, and contrastive learning. These are the foundational elements of the research problem and serve as the motivation for our work.
2.1 Problem Formulation
In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting:
$\mathcal{X}=(x_{1},x_{2},\cdots,x_{T}),$ (1)
where $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines. The objective is to train a model using $\mathcal{X}_{train}$ and provide predictions $\mathcal{Y}_{test}=(y_{1},y_{2},\cdots,y_{T^{\prime}})$ on another sequence $\mathcal{X}_{test}$ of length $T^{\prime}$. Here, $y_{t}\in\{0,1\}$, where 1 signifies an anomaly at that time point and 0 indicates a normal point.
2.2 Time Series Anomaly Detection
There exist various methods proposed for time series anomaly detection, encompassing statistical approaches, classical machine learning methods, and deep learning techniques. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) Schölkopf et al. (2001) and Support Vector Data Description (SVDD) Tax and Duin (2004). Deep learning methods include DAGMM, OmniAnomaly Su et al. (2019), USAD Audibert et al. (2020), LSTM-VAE Park et al. (2018), and DeepAnT, which are trained on normal data and detect anomalies by inference using reconstruction errors. Some techniques Kim et al. (2023),Chen et al. (2022) have leveraged self-attention mechanisms and have demonstrated promising detection results. Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like Li et al. (2021a) introduced sequence variational autoencoder model based on smoothness constraints, whereas Xu et al. (2021) proposed prior and sequence associations for temporal modeling, both achieving enhanced performance.
2.3 MLP Mixer for Time Series
The MLP Mixer, initially proposed as a novel architecture for computer vision Tolstikhin et al. (2021), has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like Nie et al. (2023); Gong et al. (2023) have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF). To the best of our knowledge, there is currently an absence of MLP Mixer application in the time series anomaly detection.
2.4 Contrastive Learning
Contrastive learning has garnered considerable attention across various domains for its capacity to train models by contrasting positive pairs against negative ones. The objective is to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs. Additionally, methods like BYOL Grill et al. (2020) and SimSiam Chen and He (2021) have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as Eldele et al. (2021) and Pranavan et al. (2022) have explored the integration of contrastive learning into time series analysis. Yang et al. (2023), on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
3 Proposed Method
3.1 Overall Architecture
Figure 1 illustrates the overall structure of PatchAD. PatchAD adopts patch sizes at multiple scales. We exemplify its architecture using a single scale. PatchAD is an $L$-layer network, where each layer comprises distinct modules: MLP Mixers, Project head, and ReWeight, forming its structure. We have devised four distinct types of MLP Mixers: Channel Mixer, Inter Mixer, Intra Mixer, and MixRep Mixer, each dedicated to learning various channels, inter-patch, intra-patch relationships, and a unified representation space. The Project head serves to prevent the model from converging to trivial solutions, while the ReWeight module manages the weights across different layers.
The focal point of our design lies in the four distinct MLP Mixers. Inspired by Yang et al. (2023), we utilize various MLP Mixers to learn different representations from the input. Normal points share a common latent space across different views, while anomalies, being scarce and lacking specific patterns, struggle to share a coherent representation with normal points. Consequently, normal points exhibit minor discrepancies across different views, whereas anomalies showcase larger variations. PatchAD leverages inter-intra discrepancy to model this relationship.
3.2 Multi-scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. This enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping blocks of patches, which is called $\operatorname{Patching}(\cdot)$. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ \mathcal{X}^{C\times T}\rightarrow\mathcal{X}^{C\times N\times P},$ (2)
where $\operatorname{PE}(\cdot)$ is positional embedding.
PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. And $\operatorname{VE}(\cdot)$ is a single linear layer.
$\mathcal{X}_{inter}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times N\times D},$ (3)
$\mathcal{X}_{intra}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times P\times D}.$
3.3 Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. Figure 1(c) illustrates the MLP block we designed. It consists of two fully connected layers, a normalization layer, a GELU non-linear activation layer, and utilizes a residual connection from input to output. This can be represented as:
$\mathcal{X}=\operatorname{FC}(\operatorname{GELU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (4)
The following are the different types of MLP Mixers utilized in PatchAD:
  1. 1.
    The Channel Mixer enables PatchAD to capture correlations among different channels.
  2. 2.
    The Inter Mixer empowers PatchAD to capture inter-patch representations, facilitating the learning of global expressions.
  3. 3.
    The Intra Mixer enables PatchAD to capture intra-patch representations, facilitating the learning of local expressions.
  4. 4.
    The MixRep Mixer allows PatchAD to embed two distinct views into the same representation space.
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation.
3.4 Dual Project Constraint
To prevent the model from converging to trivial solutions, each layer’s output requires to be constrained by Dual Project Head. Prior research suggests that employing a similar structure effectively prevents models from converging to trivial solutions. For instance, SimCLR Chen et al. (2020) uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training. The projection head comprises two layers of MLP connections, excluding non-linear and normalization layers. Hence, for both the inter-patch and intra-patch encoders, we can derive the projected representations $\mathcal{N}^{\prime}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}^{\prime}\in\mathbb{R}^{C\times P\times D}$. This computation can be represented as:
$\mathcal{N}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{N})}),\ \mathcal{P}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{P})}).$ (5)
To amalgamate outputs from different layers and prevent model degradation, we employ a simple ReWeight module to assign distinct weights to various layers. Considering representations from $L$ layers, $\{\mathcal{N}_{1},\mathcal{N}_{2},\cdots,\mathcal{N}_{L}\}$ and $\{\mathcal{P}_{1},\mathcal{P}_{2},\cdots,\mathcal{P}_{L}\}$, we can derive the final result:
$\mathcal{N}_{l}$ $=\alpha_{l}\cdot\mathcal{N}_{l},\ \alpha_{l}=\operatorname{Softmax}(\alpha),$ (6)
$\mathcal{P}_{l}$ $=\beta_{l}\cdot\mathcal{P}_{l},\ \beta_{l}=\operatorname{Softmax}(\beta).$
The introduction of ReWeight allows for a better balance between outputs from different layers, resulting in improved performance.
3.5 Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both.
The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))$ (7)
$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$
$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))$ (8)
$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
where $\operatorname{KL}(\cdot||\cdot)$ represents the KL divergence distance, and $\operatorname{StopGrad}$ denotes gradient stoppage, employed for asynchronously optimizing the two branches. The combined loss is defined as:
$\mathcal{L}_{cont}=\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}.$ (9)
Due to the mismatch in dimensions between $\mathcal{N}$ and $\mathcal{P}$, an initial upsampling of both is required for comparability. For the inter-patch view, where only differences between patches exist, we replicate the final $\mathcal{N}$ within patches. Conversely, for the intra-patch view, exhibiting anomalies between points within patches, we replicate it multiple times to obtain the final $\mathcal{P}$. This also addresses the necessity for introducing multi-scale, as it compensates for the information loss during upsampling.
To prevent model degradation, we augment the original loss function with an additional constraint from the projection head, defined as:
$\mathcal{L}_{proj}=\frac{\mathcal{L}_{\mathcal{N}^{\prime}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}+\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}^{\prime}}}{len(\mathcal{N})}.$ (10)
The final PatchAD loss consists of the aforementioned two components, with a constraint coefficient used to control the strength of the constraint:
$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}.$ (11)
3.6 Anomaly Score
Based on the assumption that normal points across different views share hidden patterns, while anomalies exhibit larger discrepancies, the difference between the two views can be employed as an anomaly score. The final anomaly score can be expressed as:
$\operatorname{AnomalyScore}(\mathcal{X})=\sum\operatorname{KL}(\mathcal{N},\mathcal{P})+\operatorname{KL}(\mathcal{P},\mathcal{N}).$ (12)
Based on the aforementioned anomaly score, we set a hyperparameter $\sigma$ to determine whether a point is considered an anomaly. If it surpasses the threshold, it is identified as an outlier.
${y}_{i}=\left\{\begin{aligned} &1\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})\geq\sigma\\ &0\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})<\sigma.\end{aligned}\right.$ (13)
Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset MSL SMAP PSM SWaT Average F1
Metric P R F1 P R F1 P R F1 P R F1
OCSVM Schölkopf et al. (2001) 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
IForest Liu et al. (2008) 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
LOF Breunig et al. (2000) 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
U-Time Perslev et al. (2019) 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
ITAD Shin et al. (2020) 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
DAGMM Zong et al. (2018) 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
VAR Anderson (1976) 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD Yairi et al. (2017) 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA Tariq et al. (2019) 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 Deldari et al. (2021) 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
BeatGAN Zhou et al. (2019) 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM-VAE Park et al. (2018) 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BOCPD Adams and MacKay (2007) 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
Deep-SVDD Ruff et al. (2018) 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
LSTM Hundman et al. (2018) 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly Su et al. (2019) 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion Li et al. (2021b) 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC Shen et al. (2020) 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans Xu et al. (2021) 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetector Yang et al. (2023) 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
PatchAD (Ours) 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
Table 2: Overall results on NIPS-TS datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
Metric P R F1 P R F1
MatrixProfile Yeh et al. (2016) 16.70 17.50 17.10 4.60 18.50 7.40 12.25
GBRT Elsayed et al. (2021) 44.70 37.50 40.80 17.50 14.00 15.60 28.20
LSTM-RNN Bontemps et al. (2016) 52.70 22.10 31.20 34.30 27.50 30.50 30.85
Autoregression Rousseeuw and Leroy (2005) 42.10 35.40 38.50 39.20 31.40 34.90 36.70
OCSVM Schölkopf et al. (2001) 47.40 49.80 48.50 18.50 74.30 29.60 39.05
IForest* Liu et al. (2008) 40.60 42.50 41.60 39.20 31.50 39.00 40.30
AutoEncoder Sakurada and Yairi (2014) 49.70 52.20 50.90 42.40 34.00 37.70 44.30
AnomalyTrans Xu et al. (2021) 90.70 47.40 62.30 25.70 28.50 27.00 44.65
IForest Liu et al. (2008) 56.90 59.80 58.30 43.90 35.30 39.10 48.70
AnomalyTrans Xu et al. (2021) 90.71 47.43 62.29 29.96 48.63 37.08 49.69
DCdetector Yang et al. (2023) 96.57 59.08 73.31 38.41 59.73 46.76 60.04
PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
4 Experiments
4.1 Setups and Evaluation Metrics
We applied eight datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others: github.com/DAMO-DI-ML/KDD2023-DCdetector. Details about these datasets are outlined in Table 4 in Appendix. Also, implementation and pseudo-code Algorithm 1 details in Appendix.
We compared our model with 27 benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; See details in Table 1 and Table 2, or in Appendix Baslines.
Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as accuracy, precision, recall, and F1-score Xu et al. (2021). Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall Huet et al. (2022), along with Volume under the surface (VUS) Paparrizos et al. (2022). Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events. Therefore, we introduce the following two new metrics. Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels. VUS considers the inclusion of anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve. Employing various methodologies allows for a more comprehensive evaluation of our model. To fairly compare with other methods, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
(a) Project constraint
(b) Anomaly ratio
(c) $d_{model}$ size
(d) Encoder layer
(e) Window size
(f) Multi-scale size
Figure 2: Parameter sensitivity studies of main hyper-parameters in PatchAD.
4.2 Comparison Results
Data Source Clarification: Due to substantial device variations, to ensure fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes. The results of AnomalyTrans and DCdetector in Table 1 and Table 2 are our reproductions.
From Table 1, PatchAD achieves the best average-level F1 performance. As far as we know, there is ongoing controversy regarding the evaluation of time series anomaly detection. However, precision, recall, and F1-score remain the most widely used evaluation metrics. To comprehensively assess the performance of our approach, additional metrics such as affiliation precision/recall and VUS have been introduced.
In Table 2, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types. PatchAD consistently achieves SOTA results on these two datasets compared to other baseline methods. Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset.
Given the remarkable performance of AnomalyTrans and DCdetector surpassing other baselines significantly, we conducted a comprehensive comparison of PatchAD with them across all datasets using multiple metrics, as depicted in Table 5 in Appendix. Our proposed method outperforms or at least remains competitive compared to these two SOTA methods. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive.
(a) Global point
(b) Contextual point
(c) Seasonal
(d) Group point
(e) Trend
Figure 3: Visualization of generated synthetic data, PatchAD anomaly scores and ground truth.
(a) PatchAD’s loss analysis
(b) Time cost comparison
(c) Time cost in different $d_{model}$
(d) Time cost in different window
Figure 4: The analysis of PatchAD and DCdetector.
4.3 Model Analysis
4.3.1 Ablation Studies
To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3. The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes the removal of the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modeling inter-channel information and sharing channel information are beneficial for model learning. Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, ’wo_constraint’ showcases the impact of removing the constraint on the projection head. In conjunction with Figure 3(a), we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
Table 3: Ablation experiments on MSL, SMAP, WADI datasets.
Dataset MSL SMAP WADI
Methods Acc P R F1 Acc P R F1 Acc P R F1
wo_ch_sharing 98.78 91.93 96.90 94.35 98.84 93.50 97.75 95.58 98.71 84.99 94.25 89.38
wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.50 97.10 95.27 98.92 85.40 97.93 91.24
wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74
wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53
Ours 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
4.3.2 Parameter Sensitivity
We conducted a sensitivity analysis on PatchAD’s parameters. The default parameter settings remain consistent with the basic configuration. Figure 1(a) demonstrates the impact of the projection head constraint on PatchAD. Overall, it holds more significance for smaller-scale datasets like MSL. However, excessively strong constraints hinder the model’s learning process. Therefore, setting this constraint between 0.1 to 0.4 is more advisable. Figure 1(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2. The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets compared to the MSL dataset. For most datasets, setting this value between 0.8 and 1 yields optimal results. Figure 1(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$. PatchAD demonstrates better results with smaller $d_{model}$ sizes. Therefore, to strike a balance between model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively. Figure 1(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4. Figure 1(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210. The time window is a critical parameter, typically set to an initial value of 105 across most datasets. Lastly, Figure 1(f) displays PatchAD’s performance under different multi-scale patch size combinations. The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
4.3.3 Visual Analysis
We visualized PatchAD’s response to various anomalies in Figure 3. Synthetic univariate time series were generated by Lai et al. (2021), incorporating anomalies such as global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies. The first row depicts the original data, while the second row displays PatchAD’s output of anomaly scores against the anomaly threshold. In particular, PatchAD has the ability to detect various types of anomalies.
4.3.4 Time Cost
We conducted a time consumption analysis on the SMD dataset under identical experimental conditions. Figure 3(b) demonstrates that under different sliding step conditions, our method outperforms DCdetector. Specifically, when the step is set to 1, PatchAD’s speed is 1.45x faster than DCdetector. Additionally, Figure 3(d) illustrates the time required for model training with a fixed window size (window size=105). This indicates that PatchAD exhibits higher efficiency in training compared to DCdetector. Figure 3(c) shows the time required for model training with a fixed sliding step (step=35) while varying the window size from 35 to 175 for different $d_{model}$ sizes. Both graphs show that the time consumption of PatchAD exhibits a linear growth with respect to the independent variables. Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD. However, the time step noticeably affects the duration as this parameter leads to increased data volume.
5 Conclusion
We presents PatchAD, an innovative algorithm tailored for time series anomaly detection. PatchAD developed a multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning. This architecture is unique in its use of a purely MLP-based approach to discern both inter-patch and intra-patch relationships in time series sequences. By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns. Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation. Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation. The results demonstrate that proposed PatchAD outperforms other methods across eight benchmark datasets.
Appendix
This is the appendix of PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection.
Datasets
Many publicly available real-world datasets are extensively used in existing methods. We utilize eight datasets to evaluate model performance: (1) MSL (Mars Science Laboratory dataset), (2) SMAP (Soil Moisture Active Passive dataset), (3) PSM (Pooled Server Metrics), (4) SMD (Server Machine Dataset), (5) SWaT (Secure Water Treatment Testbed), (6) WADI (Water Distribution Testbed), (7) NIPS-TS-SWAN, and (8) NIPSTS-GECCO. Table 4 shows the details.
Table 4: Details of benchmark datasets.
Dataset #Training #Test (Labeled) Dimension Anomaly ratio (%)
SML 58317 73729 55 10.5
SMAP 135183 427617 25 12.8
PSM 132481 87841 25 27.8
SMD 708377 708393 38 4.2
SWaT 99000 89984 26 12.2
WADI 1048571 172801 123 5.99
NIPS-TS-SWAN 60000 60000 38 32.6
NIPS-TS-GECCO 69260 69261 9 1.1
Baslines
We compared our model against 27 benchmark models, including: (1) Reconstruction-based: AutoEncoder Sakurada and Yairi [2014], LSTM-VAE Park et al. [2018], OmniAnomaly Su et al. [2019], BeatGAN Zhou et al. [2019], InterFusion Li et al. [2021b]; (2) Autoregression-based: VAR Anderson [1976], Autoregression Rousseeuw and Leroy [2005], LSTM-RNN Bontemps et al. [2016], LSTM Hundman et al. [2018], CL-MPPCA Tariq et al. [2019]; (3) Density-based: LOF Breunig et al. [2000], MMPCACD Yairi et al. [2017], DAGMM Zong et al. [2018]; (4) Clustering-based: DeepSVDD Ruff et al. [2018], THOC Shen et al. [2020], ITAD Shin et al. [2020]; Matrix Profile Yeh et al. [2016]; (5) The classic methods: OCSVM Schölkopf et al. [2001], OCSVM-based subsequence clustering (OCSVM*), IForest Liu et al. [2008], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) Elsayed et al. [2021]; (6) Change point detection and time series segmentation methods: BOCPD Adams and MacKay [2007], U-Time Perslev et al. [2019], TS-CP2 Deldari et al. [2021]; (7) Transformer-based: Anomaly Transformer  [^7] https://github.com/thuml/Anomaly-Transformer Xu et al. [2021]; DCdetector  [^8] https://github.com/DAMO-DI-ML/KDD2023-DCdetector Yang et al. [2023]; These two are current state-of-the-art methods.
Implementation Details
All experiments were conducted using PyTorch implementation on a single NVIDIA GTX 1080TI GPU. We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3). The hidden layer dimension $d_{{model}}$ is set to 40. The default projector constraint coefficient is $0.2$. The Adam optimizer’s default learning rate is $10^{-4}$, with a batch size of 128, trained for 3 epochs across all datasets. Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively. If an anomaly score at any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly. The default value for $\sigma$ is 1. For details on parameter selection, please refer to the Appendix.
Algorithm’s Pseudo-code
Algorithm 1 PatchAD’s Python-like code
PE = PositionEmbedding(d_model) # Postional Embedding
VE_n, VE_p = Linear(dim1, d_model), Linear(dim2, d_model)
# Value Embedding of N and P views
# Patching and Embedding
x = PE(x)
x_n = rearange(x, ’b (n p) c -> b c n p’, p=patch_size)
x_p = rearange(x, ’b (p n) c -> b c p n’, p=patch_size)
x_n = VE_n(x_n)
x_p = VE_p(x_p)
# x_n’s shape (B, c_dim, n_dim, d_model)
# x_p’s shape (B, c_dim, p_dim, d_model)
Channel_Mixer = MLP(c_dim, transpose_dim=1)
Inter_Mixer = MLP(n_dim, transpose_dim=2)
Intra_Mixer = MLP(p_dim, transpose_dim=2)
MixRep_Mixer = MLP(d_model, transpose_dim=3)
# Patch Mixer Encoding
x_inter = Channel_Mixer(x_n)
x_intra = Channel_Mixer(x_p)
x_inter = Inter_Mixer(x_inter)
x_intra = Intra_Mixer(x_intra)
x_inter = MixRep_Mixer(x_inter)
x_intra = MixRep_Mixer(x_intra)
# Dual Project
x_inter = x_inter.mean(1)
x_intra = x_intra.mean(1)
Inter_Proj = MLP(d_model)
Intra_Proj = MLP(d_model)
x_inter2 = Inter_Proj(x_inter)
x_intra2 = Intra_Proj(x_intra)
# Ouput shape (B, n_dim, d_model) and (B, p_dim, d_model)
# Upsample to (B, L, d_model)
# Update
def loss_fn(a, b):
    return KL(a, b.detach()) + KL(b.detach(), a)
def inter_intra_loss_fn(n, p):
    return loss_fn(n, p) - loss_fn(p, n)
loss1 = inter_intra_loss_fn(x_inter, x_intra)
loss_n_proj = inter_intra_loss_fn(x_inter, x_intra2)
loss_p_proj = inter_intra_loss_fn(x_inter2, x_intra)
loss_proj = loss_n_proj + loss_p_proj
final_loss = (1-c) * loss1 + c * loss_proj
# backward and optimize
Full Comparison Results
In Table 5, Aff-P and Aff-R are the precision and recall of affiliation metric Huet et al. [2022], respectively. R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR Paparrizos et al. [2022], respectively. V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve Paparrizos et al. [2022], respectively. Besides, we give more results in table 6, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness as well as performance.
Table 5: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset Method Acc F1 Aff-P Aff-R R_A_R R_A_P V_ROC V_PR
MSL AnomalyTrans 98.69 93.93 51.76 95.98 90.04 87.87 88.20 86.26
DCdetector 98.83 94.77 51.01 96.74 91.29 89.12 89.80 87.81
PatchAD 98.92 95.02 52.06 96.44 90.89 88.66 90.09 87.99
SMAP AnomalyTrans 99.05 96.41 51.39 98.68 96.32 94.07 95.52 93.37
DCdetector 99.06 96.38 50.29 98.04 93.76 92.40 92.70 91.47
PatchAD 99.15 96.75 52.30 98.70 96.05 94.23 95.12 93.42
PSM AnomalyTrans 98.68 97.37 55.35 80.28 91.83 93.03 88.71 90.71
DCdetector 98.81 97.63 52.33 76.95 89.91 91.74 85.57 88.52
PatchAD 98.95 98.11 55.58 82.72 94.85 95.65 89.26 91.67
SMD AnomalyTrans 99.11 90.06 72.63 98.27 82.12 75.90 81.86 75.69
DCdetector 98.78 86.02 51.61 93.44 76.13 69.85 73.60 67.45
PatchAD 98.84 86.62 50.81 93.32 77.55 71.16 74.09 67.88
SWaT AnomalyTrans 98.62 94.62 53.03 98.08 97.89 93.47 97.92 93.49
DCdetector 93.25 96.51 52.76 97.33 98.23 95.54 98.25 95.56
PatchAD 99.12 96.52 52.78 98.17 98.23 95.55 98.26 95.57
WADI AnomalyTrans 98.17 85.47 87.84 99.75 87.20 78.50 87.36 78.63
DCdetector 98.70 89.26 59.63 90.98 87.56 81.62 87.16 81.23
PatchAD 99.06 92.43 60.37 91.76 89.85 84.18 90.68 84.99
NIPS-TS-SWAN AnomalyTrans 84.57 62.29 58.45 9.49 86.42 93.26 84.81 92.00
DCdetector 85.99 73.31 38.73 3.13 88.02 94.78 86.02 93.49
PatchAD 86.05 73.41 51.44 4.34 88.03 94.80 86.13 93.57
NIPS-TS-GECCO AnomalyTrans 98.26 37.08 55.65 89.12 60.74 28.17 60.48 28.02
DCdetector 98.57 46.76 49.66 87.68 61.01 32.13 60.77 31.90
PatchAD 98.64 49.35 53.04 89.48 63.59 35.92 63.09 35.43
Table 6: More results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset MSL SMAP PSM SMD SWaT Average F1
Metric P R F1 P R F1 P R F1 P R F1 P R F1
LOF 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 56.34 39.86 46.68 72.15 65.43 68.62 60.94
OCSVM 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 44.34 76.72 56.19 45.39 49.22 47.23 60.25
U-Time 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 65.95 74.75 70.07 46.20 87.94 60.58 65.62
IForest 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 42.31 73.29 53.64 49.29 44.95 47.02 61.22
DAGMM 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 67.30 49.89 57.30 89.92 57.84 70.40 70.18
ITAD 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 86.22 73.71 79.48 63.13 52.08 57.08 70.92
VAR 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 78.35 70.26 74.08 81.59 60.29 69.34 74.66
MMPCACD 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 71.20 79.28 75.02 82.52 68.29 74.73 75.74
CL-MPPCA 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 82.36 76.07 79.09 76.78 81.50 79.07 76.66
TS-CP2 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 87.42 66.25 75.38 81.23 74.10 77.50 79.00
Deep-SVDD 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 78.54 79.67 79.10 80.42 84.45 82.39 80.97
BOCPD 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 70.90 82.04 76.07 89.46 70.75 79.01 80.33
LSTM-VAE 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 75.76 90.08 82.30 76.00 89.50 82.20 81.24
BeatGAN 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 72.90 84.09 78.10 64.01 87.46 73.92 80.24
LSTM 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 78.55 85.28 81.78 86.15 83.27 84.69 83.32
OmniAnomaly 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 83.68 86.82 85.22 81.42 84.30 82.83 84.69
InterFusion 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 87.02 85.43 86.22 80.59 85.58 83.01 85.70
THOC 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 79.76 90.95 84.99 83.94 86.36 85.13 88.01
AnomalyTrans 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 84.14 96.88 90.06 89.79 100.0 94.62 94.48
DCdetector 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 83.04 89.22 86.02 93.25 100.0 96.51 94.26
PatchAD 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 83.06 90.51 86.62 93.28 100.0 96.52 94.60
Table 7: Parameter studies on project head constraint results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
Contraint Acc P R F1 Acc P R F1 Acc P R F1
0.0 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 97.42 98.03 97.72
0.1 98.92 91.96 98.33 95.03 99.04 94.43 98.33 96.34 98.72 97.35 98.04 97.70
0.2 98.92 92.05 98.20 95.02 99.11 94.42 98.91 96.61 98.91 97.43 98.66 98.04
0.3 98.89 92.03 97.93 94.88 99.14 94.41 99.11 96.71 98.89 97.42 98.59 98.00
0.4 98.86 92.09 97.54 94.73 99.12 94.45 98.97 96.66 98.71 97.42 97.94 97.68
0.5 98.96 92.12 98.60 95.25 99.02 94.40 98.17 96.25 98.73 97.42 98.03 97.72
0.6 98.92 92.11 98.20 95.06 99.01 94.39 98.07 96.19 98.92 97.45 98.70 98.07
0.7 98.84 91.91 97.59 94.67 99.07 94.43 98.55 96.44 98.76 97.39 98.14 97.76
0.8 98.91 92.01 98.20 95.00 99.10 94.40 98.81 96.55 98.92 97.42 98.73 98.07
0.9 98.78 91.99 96.90 94.38 98.96 94.37 97.70 96.01 98.92 97.47 98.64 98.05
Table 8: Parameter studies on anomaly threshold $\sigma$ results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
AR Acc P R F1 Acc P R F1 Acc P R F1
0.5 97.93 95.39 84.39 89.55 99.15 96.46 96.91 96.68 98.45 98.61 95.77 97.17
0.6 98.27 94.70 88.54 91.52 99.15 95.84 97.58 96.70 98.77 98.33 97.20 97.76
0.7 98.17 93.76 88.54 91.07 99.09 95.22 97.77 96.48 98.70 98.07 97.20 97.64
0.8 98.45 93.19 92.04 92.61 99.15 94.73 98.86 96.75 98.95 97.94 98.29 98.11
0.9 98.45 92.36 92.93 92.64 99.01 94.07 98.49 96.23 98.92 97.68 98.45 98.06
1.0 98.84 91.80 97.73 94.67 99.01 94.05 98.49 96.22 98.70 97.30 98.03 97.66
1.1 98.73 91.22 97.36 94.19 98.85 92.76 98.69 95.63 98.68 97.06 98.20 97.63
1.2 98.64 90.45 97.36 93.78 98.76 92.19 98.69 95.33 98.62 96.85 98.20 97.52
Table 9: Parameter studies on $d_{model}$ size results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
$d_{model}$ Acc P R F1 Acc P R F1 Acc P R F1
10 98.62 91.81 95.42 93.58 99.04 94.42 98.28 96.31 98.87 97.38 98.56 97.97
20 98.92 92.08 98.20 95.04 98.98 94.33 97.88 96.07 98.73 97.37 98.05 97.71
30 98.89 92.06 97.94 94.91 99.10 94.43 98.80 96.57 98.87 97.43 98.52 97.97
40 98.91 92.03 98.20 95.02 99.13 94.49 98.95 96.67 98.73 97.42 98.03 97.72
50 98.93 92.15 98.20 95.08 99.11 94.45 98.89 96.62 98.90 97.44 98.63 98.03
60 98.91 92.03 98.20 95.02 99.13 94.44 98.98 96.66 98.94 97.48 98.74 98.10
70 98.87 92.08 97.67 94.79 99.05 94.40 98.43 96.37 98.97 97.48 98.83 98.15
80 98.77 91.99 96.70 94.29 99.03 94.38 98.24 96.27 98.89 97.44 98.58 98.00
90 98.92 92.06 98.20 95.03 99.05 94.38 98.42 96.35 98.89 97.46 98.56 98.01
Table 10: Parameter studies on encoder layers $L$ results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
Encoder layer Acc P R F1 Acc P R F1 Acc P R F1
1 98.60 91.72 95.27 93.46 98.76 94.29 96.11 95.19 98.10 97.28 95.82 96.54
2 98.88 91.96 97.93 94.85 99.01 94.42 98.09 96.22 98.79 97.44 98.20 97.82
3 98.92 92.07 98.20 95.03 99.10 94.41 98.79 96.55 98.84 97.47 98.36 97.91
4 98.88 92.19 97.67 94.85 99.11 94.42 98.91 96.61 98.95 97.41 98.83 98.12
5 98.83 91.95 97.41 94.60 99.13 94.46 99.02 96.69 98.87 97.44 98.52 97.98
Table 11: Parameter studies on window size results (window size=105). All results are in %.
Dataset MSL SMAP PSM
Window size Acc P R F1 Acc P R F1 Acc P R F1
30 98.11 91.47 90.52 90.99 98.91 94.38 97.23 95.78 98.88 97.48 98.51 97.99
45 97.98 91.23 89.44 90.33 98.93 95.55 96.14 95.84 98.67 97.30 97.92 97.61
60 98.33 92.05 92.16 92.10 98.86 94.18 97.09 95.62 98.75 97.59 97.93 97.76
75 98.51 91.79 94.31 93.03 98.93 94.27 97.60 95.90 98.72 97.35 98.06 97.70
90 98.71 91.92 96.24 94.03 99.06 94.34 98.55 96.40 98.92 97.44 98.71 98.07
105 98.92 92.05 98.20 95.02 99.08 94.44 98.60 96.48 98.66 97.38 97.79 97.59
120 98.86 92.10 97.55 94.75 99.13 94.47 98.98 96.67 98.79 97.53 98.13 97.83
135 98.65 91.85 95.67 93.72 99.07 94.46 98.50 96.44 98.88 97.38 98.60 97.99
150 98.71 91.83 96.30 94.02 99.12 94.48 98.91 96.64 98.66 97.43 97.72 97.58
165 98.61 91.85 95.27 93.53 99.09 94.50 98.65 96.53 98.73 97.44 97.98 97.71
180 98.53 91.73 94.55 93.12 99.01 94.35 98.18 96.23 98.70 97.55 97.78 97.66
195 98.77 92.02 96.70 94.31 98.96 94.35 97.74 96.01 98.87 97.44 98.50 97.97
210 98.81 91.97 97.17 94.50 98.83 94.30 96.71 95.49 98.92 97.45 98.69 98.06
Table 12: Parameter studies on multi-scale patch size results (window size=105). All results are in %.
Dataset MSL SMAP PSM
Patch size Acc P R F1 Acc P R F1 Acc P R F1
[1] 98.49 91.73 94.12 92.91 99.09 94.47 98.69 96.53 98.66 97.21 97.99 97.60
[3] 98.57 91.67 95.08 93.34 99.08 94.44 98.61 96.48 98.84 97.56 98.26 97.91
[5] 98.34 91.70 92.66 92.18 99.02 94.40 98.20 96.26 98.83 97.44 98.37 97.91
[7] 98.45 91.96 93.41 92.68 99.13 94.46 99.04 96.70 98.88 97.43 98.56 97.99
[1, 3] 98.63 91.87 95.42 93.61 99.11 94.48 98.86 96.62 98.78 97.46 98.16 97.81
[1, 5] 98.63 91.88 95.42 93.61 99.07 94.43 98.58 96.46 98.72 97.37 98.04 97.70
[1, 7] 98.63 91.89 95.42 93.62 99.04 94.43 98.32 96.33 98.71 97.39 97.99 97.69
[3, 5] 98.78 91.99 96.90 94.38 99.07 94.39 98.63 96.46 98.91 97.43 98.66 98.04
[3, 7] 98.68 91.92 95.93 93.88 99.09 94.40 98.76 96.53 98.87 97.41 98.55 97.97
[5, 7] 98.61 91.86 95.27 93.53 99.08 94.44 98.63 96.49 98.85 97.40 98.46 97.93
[1, 3, 5] 98.85 92.01 97.54 94.69 99.12 94.47 98.92 96.64 98.82 97.37 98.39 97.88
[1, 3, 7] 98.74 92.00 96.46 94.18 99.09 94.40 98.71 96.50 98.79 97.43 98.21 97.82
[1, 5, 7] 98.92 92.05 98.20 95.02 99.04 94.38 98.36 96.33 98.67 97.36 97.87 97.61
[3, 5, 7] 98.56 91.76 94.89 93.30 99.08 94.42 98.61 96.46 98.87 97.45 98.52 97.98
[1, 3, 5, 7] 98.62 91.78 95.42 93.56 99.08 94.43 98.61 96.48 98.92 97.43 98.71 98.06