### v5

# PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection

Abstract
Anomaly detection in time series analysis is a pivotal task, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios. While prior studies have largely employed reconstruction-based approaches, which limits the models’ representational capacities. Moreover, existing deep learning-based methods are not sufficiently lightweight. Addressing these issues, we present PatchAD, our novel, highly efficient multiscale patch-based MLP-Mixer architecture that utilizes contrastive learning for representation extraction and anomaly detection. With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 3.2MB. Its efficacy is demonstrated by state-of-the-art results across 9 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms. PatchAD significantly improves the classical F1 score by 50.5%, the Aff-F1 score by 7.8%, and the AUC by 10.0%. The code is publicly available.  [^1] https://github.com/EmorZz1G/PatchAD
1 Introduction
Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from expected behaviour [1, 2, 3]. With the rapid progression of large-scale sensing technologies and storage capabilities, temporal anomaly detection has found widespread application in various sectors. For instance, the Internet of Things (IoTs) serve to monitor abnormal events in sensor data [4, 5]. Unsupervised algorithms, capable of operating without labels, have gained wide popularity. These include reconstruction-based[6, 7], density-based[8, 9], contrastive learning[10, 11], autoregression-based[12, 13, 14], large language models (LLMs)[15] etc. However, extracting anomalies from extensive and complex temporal data presents significant challenges for different algorithms. Challenge 1: Accurately defining anomaly representations is challenging due to real-world anomalies’ diverse and dynamic nature. Effective anomaly detection requires models that can learn the interrelationships between various data channels. Challenge 2: There is still significant potential for improvement in terms of training and inference speed, model complexity, and the lack of lightweight features. Challenge 3: In an unsupervised setting, the challenge lies in establishing a reliable basis to model normal and anomalous data. Specifically, several algorithms have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures and model degradation issue of contrastive learning. Additionally, they either do not explicitly model inter-channel dependencies [10] or rely solely on basic MLPs or Transformers that focus on temporal information [16], mainly considering point-level features [17].
On the other hand, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points. This difference aids our model in detecting anomalies, ensuring that normal data maintains consistent representations across different views, whereas anomalies do not. This fundamental principle forms the basis of our proposed anomaly detection method.
In response to the above challenges and principles, we introduce PatchAD, a lightweight Patch-based MLP architecture for time series Anomaly Detection. Specifically, proposed Multi-scale Patching and Embedding decomposes the input data along the temporal dimension into non-overlapping patches, extracting semantically richer features beyond simple point-level characteristics. Besides, various submodels address various time scales, each focusing on different patch sizes. Patch Mixer Encoders are introduced to learn the complex relationships within and between patches and channels. A shared MixRep Mixer uniformly represents the feature space, while a Dual Project Constraint prevents the model from adopting overly simplistic solutions, thereby enhancing its representational capabilities. In detail, this work makes the following contributions:
Figure 1: Overall performances under different metrics.
Figure 2: Model sizes.
(1) State-of-the-art Anomaly Detection (Figure2): We propose a novel contrastive learning-based time series anomaly detection method. PatchAD achieves superior performance on 9 benchmark datasets, surpassing over 30 baselines across various evaluation metrics, with and without prior association.
(2) Lightweight and Efficient Design (Figure2): PatchAD employs a multi-scale patch-based MLP-Mixer architecture with different novel strategies, achieving remarkable efficiency and compactness compared to Transformer-based models.
(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by a theoretical analysis that elucidates the model’s underlying principles and design choices.
2 Preliminaries
2.1 Problem Formulation
In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting: $\mathcal{X}=(x_{1},x_{2},\cdots,x_{T})$, $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines. The objective is to train a model using $\mathcal{X}_{train}$ and provide predictions $\mathcal{Y}_{test}=(y_{1},y_{2},\cdots,y_{T^{\prime}})$ on another sequence $\mathcal{X}_{test}$ of length $T^{\prime}$. Here, $y_{t}\in\{0,1\}$, where 1 signifies an anomaly at that time point and 0 indicates a normal point.
2.2 Time Series Anomaly Detection
Figure 3: The workflow of the proposed PatchAD framework.
The time series anomaly detection methods encompass statistical approaches, classical machine learning methods, and deep learning techniques[18]. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) [19] and Support Vector Data Description (SVDD) [20]. Deep learning methods include DAGMM, OmniAnomaly [7], USAD [17], LSTM-VAE [6], and DeepAnT [21], which are trained on normal data and detect anomalies by inference using reconstruction errors. Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [22, 23]. While [24] designed six different anomaly injection methods to train the model in a self-supervised approach, [25] utilized outlier exposure and mixup to construct a self-supervised task. Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like [26] introduced sequence variational autoencoder model based on smoothness constraints, whereas [27] proposed prior and sequence associations for temporal modeling, both achieving enhanced performance. Recently approaches based on diffusion models and Large language model have also been proposed [28, 15].
2.3 MLP Mixer & Contrastive Learning for Time Series Analysis
The MLP Mixer initially proposed as a novel architecture for computer vision [29], has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like [30, 31, 32] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF).
Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs. Additionally, methods like BYOL [33] and SimSiam [34] have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as [11] and [35] have explored integrating contrastive learning into time series analysis. [10], on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
3 Proposed Method
3.1 Overall Architecture
Figure 3 illustrates the overall structure of PatchAD, which adopts patch sizes at multiple scales. We exemplify its architecture using a single scale. PatchAD is an $L$-layer network, where each layer comprises distinct modules: MLP Mixers, Project head, and ReWeight, forming its structure. We have devised four distinct types of MLP Mixers: Channel Mixer, Inter Mixer, Intra Mixer, and MixRep Mixer, each dedicated to learning various channels, inter-patch, intra-patch relationships, and a unified representation space. The Project head serves to prevent the model from converging to trivial solutions, while the ReWeight module manages the weights across different layers.
The focal point of our design lies in the four distinct MLP Mixers. We utilize various MLP Mixers to learn different representations from the input. Normal points share a common latent space across different views, while anomalies, being scarce and lacking specific patterns, struggle to share a coherent representation with normal points. Consequently, normal points exhibit minor discrepancies across different views, whereas anomalies showcase larger variations. PatchAD leverages inter-intra discrepancy to model this relationship.
Our anomaly detection approach, PatchAD, leverages the principle that normal time series points, characterized by low information entropy, exhibit consistent representations across diverse perspectives. In contrast, anomalies with higher entropy struggle to maintain such consistency (Figure 3(c)). PatchAD employs an Inter-Patch Encoder to capture detailed features and an Intra-Patch Encoder to learn generalized trends. This combined approach enables a comprehensive understanding of normal and anomalous representations, achieving effective anomaly detection.
3.2 Multi-Scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. This enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ \mathcal{X}^{C\times T}\rightarrow\mathcal{X}^{C\times N\times P},$ (1)
where $\operatorname{PE}(\cdot)$ is positional embedding. PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. And $\operatorname{VE}(\cdot)$ is a single linear layer.
$\mathcal{X}_{inter}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times N\times D},$ (2)
$\mathcal{X}_{intra}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times P\times D}.$
3.3 Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. The MLP Mixer consists of two fully connected layers, a normalization layer and a GELU non-linear activation layer, and utilizes a residual connection from input to output. This can be represented as:
$\mathcal{X}=\operatorname{FC}(\operatorname{GELU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (3)
The following are the different types of MLP Mixers utilized in PatchAD:
  1. 1.
    Channel Mixer captures correlations among different channels.
  2. 2.
    Inter Mixer captures inter-patch representations, improving detailed expressions.
  3. 3.
    Intra Mixer captures intra-patch representations, improving generalized expressions.
  4. 4.
    MixRep Mixer embeds two distinct views into the same representation space.
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation. In the end, two separate simple MLPs are utilized to reconstruct the original data, which are then combined.
$\hat{\mathcal{X}}_{1}=\operatorname{Rec1}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec2}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (4)
3.4 Dual Project Constraint
To prevent the model from converging to trivial solutions, each layer’s output requires to be constrained by Dual Project Head. Prior research suggests that employing a similar structure effectively prevents models from converging to trivial solutions. For instance, SimCLR [36] uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training. The projection head comprises two layers of MLP connections, excluding non-linear and normalization layers. Hence, for both the inter-patch and intra-patch encoders, we can derive the projected representations $\mathcal{N}^{\prime}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}^{\prime}\in\mathbb{R}^{C\times P\times D}$. This computation can be represented as:
$\mathcal{N}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{N})}),\ \mathcal{P}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{P})}).$ (5)
To amalgamate outputs from different layers and prevent model degradation, we employ a simple ReWeight module to assign distinct weights to various layers. Considering representations from $L$ layers, $\{\mathcal{N}_{1},\mathcal{N}_{2},\cdots,\mathcal{N}_{L}\}$ and $\{\mathcal{P}_{1},\mathcal{P}_{2},\cdots,\mathcal{P}_{L}\}$, we can derive the final result:
$\mathcal{N}_{l}$ $=\alpha_{l}\cdot\mathcal{N}_{l},\ \alpha_{l}=\operatorname{Softmax}(\alpha),$ (6)
$\mathcal{P}_{l}$ $=\beta_{l}\cdot\mathcal{P}_{l},\ \beta_{l}=\operatorname{Softmax}(\beta).$
The introduction of ReWeight allows for a better balance between outputs from different layers, resulting in improved performance.
3.5 Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both. The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$ (7)
$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
where $\operatorname{KL}(\cdot||\cdot)$ represents the KL divergence distance, and $\operatorname{StopGrad}$ denotes gradient stoppage, employed for asynchronously optimizing the two branches. The combined loss is defined as:
$\mathcal{L}_{cont}=\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}.$ (8)
Due to the mismatch in dimensions between $\mathcal{N}$ and $\mathcal{P}$, an initial upsampling of both is required for comparability. For the inter-patch view, where only differences between patches exist, we replicate the final $\mathcal{N}$ within patches. Conversely, for the intra-patch view, exhibiting anomalies between points within patches, we replicate it multiple times to obtain the final $\mathcal{P}$. This also addresses the necessity for introducing multi-scale, as it compensates for the information loss during upsampling.
To prevent model degradation, we augment the original loss function with an additional constraint from the projection head, defined as:
$\mathcal{L}_{proj}=\frac{\mathcal{L}_{\mathcal{N}^{\prime}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}+\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}^{\prime}}}{len(\mathcal{N})}.$ (9)
The final PatchAD loss comprises the two aforementioned components and the reconstruction term ($\mathcal{L}_{rec}=\operatorname{MSE}(\hat{\mathcal{X}},\mathcal{X})$), which is accompanied by a constraint coefficient to regulate the strength of the constraint.
$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}+\mathcal{L}_{rec}.$ (10)
3.6 Anomaly Score
Based on the assumption that normal points across different views share hidden patterns, while anomalies exhibit larger discrepancies, the difference between the two views can be employed as an anomaly score. The final anomaly score can be expressed as:
$\operatorname{AnomalyScore}(\mathcal{X})=\sum\operatorname{KL}(\mathcal{N},\mathcal{P})+\operatorname{KL}(\mathcal{P},\mathcal{N}).$ (11)
Based on the aforementioned anomaly score, we set a hyperparameter $\sigma$ to determine whether a point is considered an anomaly. If it surpasses the threshold, it is identified as an outlier.
Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best in Bold, and the second in underlined.
Dataset MSL SMAP PSM SWaT Average F1
Metric P R F1 P R F1 P R F1 P R F1
OCSVM [19] 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
IForest [37] 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
LOF [8] 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
U-Time [38] 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
ITAD [39] 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
DAGMM [9] 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
VAR [12] 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD [40] 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA [41] 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 [42] 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
BeatGAN [43] 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM-VAE [6] 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BOCPD [44] 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
Deep-SVDD [45] 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
LSTM [46] 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly [7] 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion [47] 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC [48] 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans [27] 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetector [10] 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
TimesNet [14] 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
PatchTST [49] 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
GPT2-Adapter [15] 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
NPSR-pt [50] 10.60 100.0 19.16 24.99 96.37 39.68 27.74 100.0 43.43 12.29 100.0 21.89 31.04
NPSR-seq [50] 17.66 100.0 30.02 12.79 100.0 22.68 37.87 99.95 54.81 86.76 88.58 87.66 48.80
PatchAD (Ours) 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
4 Experiments
4.1 Setups and Evaluation Metrics
We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP (3) PSM (4) SMD (5) SWaT (6) WADI (7) NIPS-TS-SWAN (8) NIPSTS-GECCO (9) UCR. Details about these datasets are outlined in Table 4 in Appendix. Also, implementation and pseudo-code Algorithm 1 details in Appendix.
We compared our model with several benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; (8) Large language model; (9) Diffusion model; See details in Table 1 and Table 5, or in Appendix B.2.
Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as AUC, accuracy, precision, recall, F1-score with point adjustment (F1), and original F1-score (F1_ori) [27]. Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall [51], along with Volume under the surface (VUS) [52]. Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events. Therefore, we introduce the following two new metrics. Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels. VUS considers including anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve. Employing various methodologies allows for a more comprehensive evaluation of our model. To compare with other methods fairly, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
4.2 Comparison Results
Data Source Clarification: Due to substantial device variations, to ensure a fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes. The results of AnomalyTrans and DCdetector in Table 1 and Table 5 are our reproductions.
From Table 1, PatchAD achieves the best average-level F1 performance. As far as we know, there is ongoing controversy regarding the evaluation of time series anomaly detection. However, precision, recall, and F1-score remain the most widely used evaluation metrics. To comprehensively assess the performance of our approach, additional metrics such as affiliation precision/recall and VUS have been introduced.
In Table 5 in Appendix, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types. Compared to other baseline methods, PatchAD consistently achieves SOTA results on these two datasets. Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset. Given the remarkable performance of AnomalyTrans and DCdetector surpassing other baselines significantly, we conducted a comprehensive comparison of PatchAD with them across all datasets using multiple metrics, as depicted in Table 7 in Appendix. Our proposed method outperforms or at least remains competitive compared to these two SOTA methods. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive. Table 2 demonstrates that our approach outperforms previous methods on univariate dataset UCR. This shows that our method can also fetch state-of-the-art performance on univariate datasets.
Table 2: Results on UCR dataset.
Metric Acc P R F1
AnomalyTrans 99.49 60.41 100 73.08
DCdetector 99.51 61.62 100 74.05
PatchAD (Ours) 99.52 72.72 100 84.00
Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (WADI, SWAT, PSM, NIPS-TS-SWAN). PatchAD has demonstrated a remarkable improvement in the F1 score, with a notable increase of 50.5%. Additionally, there has been a 7.8% improvement in the Aff-F1 score and a commendable 10.0% boost in the AUC.
(a) Global point
(b) Contextual point
(c) Seasonal
(d) Group point
(e) Trend
Figure 4: Visualization of generated synthetic data, PatchAD anomaly scores and ground truth.
(a) Loss analysis
(b) Time cost comparison
(c) Different $d_{model}$
(d) Different window
Figure 5: The analysis of PatchAD and DCdetector.
4.3 Unveiling the Mechanisms
PatchAD leverages Inter and Intra Mixers to learn detailed and generalized features respectively, as demonstrated in Figure 8. Utilizing multiple patch sizes, as explored in Appendix (Figures 8(a) & 8(b)), enhances feature learning and generalization. The interplay between these mixers, visualized in Figure 3(c), effectively separates and amplifies distinct features, improving anomaly detection. Empirical validation, depicted in Figure 9, reveals that increasing patch size leads to more significant features acquired by the Intra Mixer, accelerating model learning. Finally, Figure 10 shows that anomalies are most easily detected in the Intra Mixer view due to its enhanced distinctiveness. In essence, maximizing the disparity between the Inter and Intra Mixers enhances PatchAD’s detection capabilities while mitigating overfitting. More details are provided in Appendix.
4.4 Model Analysis
4.4.1 Ablation Studies
Table 3: Ablation studies on MSL, SMAP, WADI datasets.
Dataset MSL SMAP WADI
Methods Acc P R F1 Acc P R F1 Acc P R F1
wo_ch_sharing 98.78 91.93 96.90 94.35 98.84 93.50 97.75 95.58 98.71 84.99 94.25 89.38
wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.50 97.10 95.27 98.92 85.40 97.93 91.24
wo_mixrep_mixer 96.41 89.54 74.66 81.43 95.25 91.52 68.55 78.70 95.76 67.55 51.99 58.76
wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74
wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53
PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3. The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes removing the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modelling inter-channel information and sharing channel information benefit model learning. Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, ’wo_constraint’ showcases the impact of removing the constraint on the projection head. In conjunction with Figure 5(a), we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
4.4.2 Parameter Sensitivity
Figure 6: Parameter sensitivity studies of main hyper-parameters in PatchAD.
A sensitivity analysis was conducted to investigate the impact of various PatchAD parameters on model performance (Figure 6). We find that a projection head constraint within the range of 0.1 to 0.4 is generally advisable, particularly for smaller datasets (Figure 6(a)). Optimal anomaly thresholds ($\sigma$) typically fall between 0.8 and 1.0, demonstrating robustness across various datasets (Figure 6(b)). A network dimension ($d_{model}$) of 40 strikes a balance between model complexity and performance (Figure 6(c)). PatchAD exhibits consistent performance across a range of window sizes (Figure 6(e)), and employing 3 or 4 encoder layers is found to be suitable (Figure 6(d)). Finally, combining multiple patch sizes demonstrates a positive impact on overall performance (Figure 6(f)). Detailed analysis for each parameter is provided in Appendix.
4.4.3 Visual Analysis
We visualized PatchAD’s response to various anomalies in Figure 4. Synthetic univariate time series were generated by [53], incorporating global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies. The first row depicts the original data, while the second row displays PatchAD’s output of anomaly scores against the anomaly threshold. In particular, PatchAD has the ability to detect various types of anomalies. Figure 7 presents the practical performance of PatchAD on two real-world datasets. PatchAD exhibits the capability to detect the majority of anomalies, albeit with two anomalies remaining undetected. Furthermore, a noticeable disparity is observed between the reconstructed features by PatchAD and the original features, indicating its proficiency in learning the underlying patterns within the data, rather than simply replicating the original features.
Figure 7: The case studies on SWAT and WADI datasets.
4.4.4 Time Cost
We conducted a time consumption analysis on the SMD dataset under identical experimental conditions. Figure 5(b) demonstrates that under different sliding step conditions, our method outperforms DCdetector. Specifically, when the step is set to 1, PatchAD’s speed is 1.45x faster than DCdetector. Additionally, Figure 5(d) illustrates the time required for model training with a fixed window size (window size=105). This indicates that PatchAD exhibits higher efficiency in training compared to DCdetector. Figure 5(c) shows the time required for model training with a fixed sliding step (step=35) while varying the window size from 35 to 175 for different $d_{model}$ sizes. Both graphs show that the time consumption of PatchAD exhibits a linear growth with respect to the independent variables. Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD. However, the time step noticeably affects the duration as this parameter increases data volume.
4.4.5 Model Size
Figure 2 compares the sizes of different models on SWAT dataset. Our model has a compact size of only 3.2MB, which is smaller than both AnomTrans and DCdetector, and significantly smaller than the LLM model (GPT2-Adapter) and diffusion model (D3R). This observation indicates that our model is more lightweight and resource-efficient.
5 Conclusion
We presents PatchAD, an innovative algorithm tailored for time series anomaly detection. PatchAD developed a lightweight multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning. This architecture is unique in using a purely MLP-based approach to discern inter-patch and intra-patch relationships in time series sequences. By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns. Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation. Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation. The results demonstrate that proposed lightweight PatchAD outperforms over 30 methods across nine benchmark datasets from different application scenarios.
Appendix
This is the appendix of PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection.
Appendix A Theory
In this section, our objective is to perform a theoretical analysis of PatchAD’s effectiveness. We want to prove that normal time series data points exhibit reduced information entropy. We have made the following assumptions based on the theoretical analysis presented in previous studies.
Assumption 1: The distribution of normal data differs from that of anomalies, and both consist of two components, denoted as $\tau+\alpha$. Here, $\tau$ represents the underlying data pattern, while $\alpha$ represents the pattern of normal or anomalous data. Therefore, we denote normal data distribution by $\Pi$ and anomalies as $\Omega$. As a result, the normal data $\mathcal{X}_{norm}$ and the anomalous data $\mathcal{X}_{anom}$ can be expressed as follows:
$\mathcal{X}_{norm}=\tau+\pi,\text{ and }\mathcal{X}_{anom}=\tau+\omega.$ (12)
Assumption 2: The number of normal data samples in the dataset is significantly larger than that of anomalous data samples. Therefore, we have the following relationship:
$|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|,$ (13)
where ||$\cdot$|| ⋅ | represents the number of samples in the data, and $\zeta$ is a large number. We denote the entire dataset as $\mathcal{D}=\{\mathcal{X}_{norm},\mathcal{X}_{anom}\}$.
Proof 1: The entropy of the anomalous data is higher than that of normal data. For a random variable $X$, its entropy is defined as, and the entropy serves as a metric for quantifying the uncertainty and complexity of the data [54]:
$\operatorname{H}(X)=-\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (14)
[…]
Using the addition rule of probability, we have $p(\tau+\pi)=p(\tau)\cdot p(\pi)$. Substituting this into the previous equation, we obtain: $\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)\log(p(\tau)\cdot p(\pi))d(\tau+\pi).$ Expanding further, we can deduce:
$\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)(\log p(\tau)+\log p(\pi))d(\tau+\pi).$ (15)
[…]
Due to the significantly larger number of samples in normal data compared to anomalous data, i.e., $|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|$, we can extract the large number $\zeta$ and use the properties of logarithm to move it inside the $\log$ function:
$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ (16)
The second term $-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ is a constant term, denoted as $C$.
Therefore, we can express the entropy of normal data as:
$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)+C.$ (17)
Similarly, we can derive the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$:
$\operatorname{H}(\mathcal{X}_{anom})=-\sum_{x\in\mathcal{X}_{anom}}P^{\prime}(x)\log P^{\prime}(x).$ (18)
At this point, we can conclude that $\operatorname{H}(\pi)<\operatorname{H}(\omega)$. Therefore, the entropy of normal data $\operatorname{H}(\mathcal{X}_{norm})$ is smaller than the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$.
Appendix B Experimental Settings
B.1 Datasets
We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others:http://github.com/DAMO-DI-ML/KDD2023-DCdetector. (9) UCR  [^7] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/UCR_TimeSeriesAnomalyDatasets2021.zip Table 4 shows the details.
Table 4: Details of benchmark datasets.
Dataset #Training #Test (Labeled) Dimension Anomaly ratio (%)
MSL 58317 73729 55 10.5
SMAP 135183 427617 25 12.8
PSM 132481 87841 25 27.8
SMD 708377 708393 38 4.2
SWaT 99000 89984 26 12.2
WADI 1048571 172801 123 5.99
NIPS-TS-SWAN 60000 60000 38 32.6
NIPS-TS-GECCO 69260 69261 9 1.1
B.2 Baslines
We compared our model against 35 benchmark models, including: (1) Reconstruction-based: AutoEncoder [55], LSTM-VAE [6], OmniAnomaly [7], BeatGAN [43], InterFusion [47], USAD [17], COUTA[24]; (2) Autoregression-based: VAR [12], Autoregression [56], LSTM-RNN [13], LSTM [46], CL-MPPCA [41], TimesNet [14], NPSR-pt and NPSR-seq [50]; (3) Density-based: LOF [8], MMPCACD [40], DAGMM [9]; (4) Clustering-based: DeepSVDD [45], THOC [48], ITAD [39]; Matrix Profile [57]; (5) The classic methods: OCSVM [19], OCSVM-based subsequence clustering (OCSVM*), IForest [37], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) [58]; (6) Change point detection and time series segmentation methods: BOCPD [44], U-Time [38], TS-CP2 [42]; (7) Transformer-based: Anomaly Transformer  [^8] https://github.com/thuml/Anomaly-Transformer [27], DCdetector  [^9] https://github.com/DAMO-DI-ML/KDD2023-DCdetector [10], TranAD [59]; (8) Large Language Model: GPT2-Adapter[15]; (9) Diffusion model: D3R [60]; These two are current state-of-the-art methods.
B.3 Implementation Details
All experiments used PyTorch implementation on a single NVIDIA GTX 1080TI GPU. We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3). The hidden layer dimension $d_{{model}}$ is set to 40. The default projector constraint coefficient is $0.2$. The Adam optimizer’s default learning rate is $10^{-4}$, with a batch size of 128, trained for 3 epochs across all datasets. Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively. If an anomaly score of any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly. The default value for $\sigma$ is 1.
Appendix C Algorithm’s Pseudo-code
Algorithm 1 PatchAD’s Python-like code
PE = PositionEmbedding(d_model) # Postional Embedding
VE_n, VE_p = Linear(dim1, d_model), Linear(dim2, d_model)
# Value Embedding of N and P views
# Patching and Embedding
x = PE(x)
x_n = rearange(x, ’b (n p) c -> b c n p’, p=patch_size)
x_p = rearange(x, ’b (p n) c -> b c p n’, p=patch_size)
x_n = VE_n(x_n)
x_p = VE_p(x_p)
# x_n’s shape (B, c_dim, n_dim, d_model)
# x_p’s shape (B, c_dim, p_dim, d_model)
Channel_Mixer = MLP(c_dim, transpose_dim=1)
Inter_Mixer = MLP(n_dim, transpose_dim=2)
Intra_Mixer = MLP(p_dim, transpose_dim=2)
MixRep_Mixer = MLP(d_model, transpose_dim=3)
# Patch Mixer Encoding
x_inter = Channel_Mixer(x_n)
x_intra = Channel_Mixer(x_p)
x_inter = Inter_Mixer(x_inter)
x_intra = Intra_Mixer(x_intra)
x_inter = MixRep_Mixer(x_inter)
x_intra = MixRep_Mixer(x_intra)
# Dual Project
x_inter = x_inter.mean(1)
x_intra = x_intra.mean(1)
Inter_Proj = MLP(d_model)
Intra_Proj = MLP(d_model)
x_inter2 = Inter_Proj(x_inter)
x_intra2 = Intra_Proj(x_intra)
# Ouput shape (B, n_dim, d_model) and (B, p_dim, d_model)
# Upsample to (B, L, d_model)
rec1 = Rec1(x_inter)
rec2 = Rec2(x_intra)
rec_x = rec1 + rec2
# Update
def loss_fn(a, b):
return KL(a, b.detach()) + KL(b.detach(), a)
def inter_intra_loss_fn(n, p):
return loss_fn(n, p) - loss_fn(p, n)
loss1 = inter_intra_loss_fn(x_inter, x_intra)
loss_n_proj = inter_intra_loss_fn(x_inter, x_intra2)
loss_p_proj = inter_intra_loss_fn(x_inter2, x_intra)
loss_proj = loss_n_proj + loss_p_proj
loss_rec = MSE(rec_x, x)
final_loss = (1-c) * loss1 + c * loss_proj + loss_rec
# backward and optimize
Appendix D Unveiling the Mechanisms
Figure 8 illustrates the knowledge acquired by PatchAD in four distinct cases. Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of $[1,2,5]$. The "Original" represents the original data, while "Rec 1" and "Rec 2" correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively. It is observed that the Intra Mixer captures the generalized knowledge in the data, which encompasses the overall trend information. Conversely, the Inter Mixer learnsthe detailed features of the data, compensating for any aspects that may not be captured by the Intra Mixer. A comparison between case 1 and case 2 reveals that using three distinct patch scales facilitates the acquisition of more detailed representations and enhances the generalization of the acquired features. Note that the y-axis ranges for Figure 8(a) and Figure 8(b) are different. Moreover, these two mixers collaborate to segregate and amplify the distinctive features effectively. The decoupling of features assists PatchAD in amplifying the anomaly-related characteristics, as indicated in Figure 3(c), thereby enhancing the detectability of anomalies. The superior performance of PatchAD can also be attributed to the varying sizes of the Inter and Intra Mixer branches, which prevent overfitting. This is the reason why PatchAD does not require the construction of negative samples in contrastive learning, as the two different networks construct implicit positive and negative samples.
Moreover, we empirically validate this conclusion through experiments. Case 1 was chosen as the baseline, and we investigated the entropy of the features acquired by the Inter and Intra Mixers under varying patch size conditions. Figures 9(a) and 9(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process. This observation suggests that the models effectively acquire knowledge and reduce uncertainty. Furthermore, we noticed that the Inter Mixer demonstrates the highest learning speed and exhibits the most substantial decrease in entropy when the patch size is configured as [1, 2, 5]. Additionally, as the patch size decreases, the entropy of the Inter Mixer increases, and it actually rises when the patch size is set to 1. This implies that the network is required to learn an excessive number of features. Figure 9(c) computes the ratio of Intra to Inter entropy, and our findings indicate that as the patch size increases, the features acquired by the Intra Mixer gain greater significance, thereby accelerating model learning through the utilization of multi-scale patch sizes.
Figure 10 illustrates the visualization of feature distributions in the SWAT dataset under three conditions. It is observed that anomalies are more easily detected in the Intra Mixer view (See Figure 10(b) ), as its distinctiveness is more pronounced compared to the other two views (Figure 10 (a) & (c) ).
According to our analysis, maximizing the disparities between the Inter Mixer and Intra Mixer can enhance the model’s detection capabilities. During network optimization, the Inter-Intra loss compels the Intra Mixer to acquire more generalized representations, effectively mitigating model overfitting.
Figure 8: Visualization of the features learnt by Inter and Intra Mixer.
Figure 9: Entropy analysis.
Figure 10: Latent distribution’s visualization by T-SNE.
Appendix E Parameter Sensitivity
We conducted a sensitivity analysis on PatchAD’s parameters. The default parameter settings remain consistent with the basic configuration. Figure 6(a) demonstrates the impact of the projection head constraint on PatchAD. Overall, it holds more significance for smaller-scale datasets like MSL. However, excessively strong constraints hinder the model’s learning process. Therefore, setting this constraint between 0.1 to 0.4 is more advisable. Figure 6(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2. The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets than the MSL dataset. Setting this value between 0.8 and 1 for most datasets yields optimal results. Figure 6(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$. PatchAD demonstrates better results with smaller $d_{model}$ sizes. Therefore, to balance model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively. Figure 6(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4. Figure 6(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210. The time window is a critical parameter, typically set to an initial value of 105 across most datasets. Lastly, Figure 6(f) displays PatchAD’s performance under different multi-scale patch size combinations. (e.g., a patch size of [3,5] represents using two scales in PatchAD.) The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
Appendix F Visualization comparison
Figure 11 shows the performance of WindowAD algorithms in the face of different types of anomalies. These methods have difficulty in detecting contextual and group point anomalies and are able to detect seasonal anomalies, but there are some false positives, which indicates that it has a low precision.
(a) Contextual
(b) Seasonal
(c) Group
Figure 11: Visualization of WindowAD-based algorithm’s score.
Appendix G Full Comparison Results
Table 6 presents the performance of different methods on four datasets, namely WADI, SWAT, PSM, and SWAN. This is the original data of Figure2. We have selected representative algorithms for experimental comparison.
Table 5: Overall results on NIPS-TS datasets. All results are in %, the best in Bold, and the second in underlined.
Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
Metric P R F1 P R F1
MatrixProfile [57] 16.70 17.50 17.10 4.60 18.50 7.40 12.25
GBRT [58] 44.70 37.50 40.80 17.50 14.00 15.60 28.20
LSTM-RNN [13] 52.70 22.10 31.20 34.30 27.50 30.50 30.85
Autoregression [56] 42.10 35.40 38.50 39.20 31.40 34.90 36.70
OCSVM [19] 47.40 49.80 48.50 18.50 74.30 29.60 39.05
IForest* [37] 40.60 42.50 41.60 39.20 31.50 39.00 40.30
AutoEncoder [55] 49.70 52.20 50.90 42.40 34.00 37.70 44.30
AnomalyTrans [27] 90.70 47.40 62.30 25.70 28.50 27.00 44.65
IForest [37] 56.90 59.80 58.30 43.90 35.30 39.10 48.70
AnomalyTrans [27] 90.71 47.43 62.29 29.96 48.63 37.08 49.69
DCdetector [10] 96.57 59.08 73.31 38.41 59.73 46.76 60.04
PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
Table 6: The performance of algorithms on different evaluation metrics without point adjustment.
Dataset WADI SWAT
Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
Deep SVDD 4.05 38.53 7.34 50.06 55.79 52.77 40.53 12.74 98.50 22.57 53.58 99.99 69.77 86.81
USAD 5.74 100.00 10.86 54.92 100.00 70.90 46.39 12.17 100.00 21.70 53.00 100.00 69.28 88.66
COUTA 98.72 15.58 26.92 99.18 23.08 37.44 49.48 94.25 6.33 11.87 79.98 33.20 46.92 75.54
TranAD 6.49 63.89 11.78 55.91 91.96 69.54 52.60 14.70 96.40 25.50 55.65 99.66 71.42 88.90
TimesNet 64.04 10.24 17.65 65.25 39.30 49.05 65.08 11.05 3.21 4.98 57.84 93.77 71.55 26.76
AnomTrans 4.14 14.12 6.40 45.73 40.66 43.05 49.20 4.81 29.32 8.27 53.03 98.08 36.08 19.20
DCdetector 5.80 14.42 8.27 59.63 90.98 69.98 49.88 15.49 5.54 8.16 52.76 97.33 68.94 50.68
GPT2-Adapter 5.68 5.50 5.59 55.62 98.29 71.04 50.01 12.16 0.96 1.78 52.51 98.13 68.41 50.00
D3R 6.99 80.99 12.86 56.47 99.98 72.17 51.39 32.87 75.97 45.89 61.47 78.52 68.96 79.95
PatchAD (Ours) 89.48(2.08) 31.22(0.74) 46.28(1.09) 73.07(0.41) 86.07(0.28) 79.04(0.12) 79.89(0.55) 85.42(0.98) 70.34(0.18) 77.15(0.48) 73.62(0.52) 84.65(0.22) 78.75(0.20) 91.19(0.12)
Dataset PSM SWAN
Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
Deep SVDD 28.59 99.91 44.46 53.92 99.98 70.06 78.09 41.74 85.03 55.99 56.84 89.93 69.66 75.32
USAD 32.36 92.17 47.90 55.30 98.71 70.88 64.53 39.73 77.94 52.63 55.26 68.62 61.22 66.76
COUTA 100.00 0.04 0.08 100.00 5.12 9.75 72.45 62.96 0.09 0.17 69.25 0.10 0.19 71.38
TranAD 35.90 54.23 43.20 60.88 92.09 73.30 65.22 38.84 78.29 51.93 57.86 90.18 70.49 70.00
TimesNet 64.85 1.27 2.49 77.84 67.23 72.15 58.74 38.63 48.89 43.16 60.75 81.29 69.53 53.62
AnomTrans 26.05 84.32 39.81 55.35 80.28 50.09 52.18 31.19 79.87 44.86 58.45 9.49 0.54 45.38
DCdetector 27.05 19.59 22.72 52.33 76.95 65.88 49.62 32.45 98.62 48.83 38.73 3.13 7.50 50.35
GPT2-Adapter 28.31 46.64 35.24 54.33 95.16 69.17 50.74 40.00 52.05 45.23 61.17 69.55 65.09 58.77
D3R 28.96 93.65 44.24 53.84 100.00 69.99 60.64 36.12 78.38 49.45 57.70 96.06 72.10 62.54
PatchAD (Ours) 47.13(0.44) 49.32(0.97) 48.20(0.64) 63.52(0.82) 96.84(1.21) 76.72(0.19) 64.74(0.01) 51.76(0.18) 65.31(0.47) 57.75(0.08) 66.48(0.80) 79.39(1.54) 72.36(0.20) 72.91(0.02)
In Table 7, Aff-P and Aff-R are the precision and recall of affiliation metric [51], respectively. R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR [52], respectively. V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve [52], respectively. Besides, we give more results in table 8, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness and performance.
Table 7: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
Dataset Method Acc F1 Aff-P Aff-R R_A_R R_A_P V_ROC V_PR
MSL AnomalyTrans 98.69 93.93 51.76 95.98 90.04 87.87 88.20 86.26
DCdetector 98.83 94.77 51.01 96.74 91.29 89.12 89.80 87.81
PatchAD 98.92 95.02 52.06 96.44 90.89 88.66 90.09 87.99
SMAP AnomalyTrans 99.05 96.41 51.39 98.68 96.32 94.07 95.52 93.37
DCdetector 99.06 96.38 50.29 98.04 93.76 92.40 92.70 91.47
PatchAD 99.15 96.75 52.30 98.70 96.05 94.23 95.12 93.42
PSM AnomalyTrans 98.68 97.37 55.35 80.28 91.83 93.03 88.71 90.71
DCdetector 98.81 97.63 52.33 76.95 89.91 91.74 85.57 88.52
PatchAD 98.95 98.11 55.58 82.72 94.85 95.65 89.26 91.67
SMD AnomalyTrans 99.11 90.06 72.63 98.27 82.12 75.90 81.86 75.69
DCdetector 98.78 86.02 51.61 93.44 76.13 69.85 73.60 67.45
PatchAD 98.84 86.62 50.81 93.32 77.55 71.16 74.09 67.88
SWaT AnomalyTrans 98.62 94.62 53.03 98.08 97.89 93.47 97.92 93.49
DCdetector 93.25 96.51 52.76 97.33 98.23 95.54 98.25 95.56
PatchAD 99.12 96.52 52.78 98.17 98.23 95.55 98.26 95.57
WADI AnomalyTrans 98.17 85.47 87.84 99.75 87.20 78.50 87.36 78.63
DCdetector 98.70 89.26 59.63 90.98 87.56 81.62 87.16 81.23
PatchAD 99.06 92.43 60.37 91.76 89.85 84.18 90.68 84.99
NIPS-TS-SWAN AnomalyTrans 84.57 62.29 58.45 9.49 86.42 93.26 84.81 92.00
DCdetector 85.99 73.31 38.73 3.13 88.02 94.78 86.02 93.49
PatchAD 86.05 73.41 51.44 4.34 88.03 94.80 86.13 93.57
NIPS-TS-GECCO AnomalyTrans 98.26 37.08 55.65 89.12 60.74 28.17 60.48 28.02
DCdetector 98.57 46.76 49.66 87.68 61.01 32.13 60.77 31.90
PatchAD 98.64 49.35 53.04 89.48 63.59 35.92 63.09 35.43
Table 8: More results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
Dataset MSL SMAP PSM SMD SWaT Average F1
Metric P R F1 P R F1 P R F1 P R F1 P R F1
LOF 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 56.34 39.86 46.68 72.15 65.43 68.62 60.94
OCSVM 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 44.34 76.72 56.19 45.39 49.22 47.23 60.25
U-Time 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 65.95 74.75 70.07 46.20 87.94 60.58 65.62
IForest 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 42.31 73.29 53.64 49.29 44.95 47.02 61.22
DAGMM 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 67.30 49.89 57.30 89.92 57.84 70.40 70.18
ITAD 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 86.22 73.71 79.48 63.13 52.08 57.08 70.92
VAR 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 78.35 70.26 74.08 81.59 60.29 69.34 74.66
MMPCACD 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 71.20 79.28 75.02 82.52 68.29 74.73 75.74
CL-MPPCA 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 82.36 76.07 79.09 76.78 81.50 79.07 76.66
TS-CP2 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 87.42 66.25 75.38 81.23 74.10 77.50 79.00
Deep-SVDD 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 78.54 79.67 79.10 80.42 84.45 82.39 80.97
BOCPD 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 70.90 82.04 76.07 89.46 70.75 79.01 80.33
LSTM-VAE 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 75.76 90.08 82.30 76.00 89.50 82.20 81.24
BeatGAN 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 72.90 84.09 78.10 64.01 87.46 73.92 80.24
LSTM 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 78.55 85.28 81.78 86.15 83.27 84.69 83.32
OmniAnomaly 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 83.68 86.82 85.22 81.42 84.30 82.83 84.69
InterFusion 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 87.02 85.43 86.22 80.59 85.58 83.01 85.70
THOC 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 79.76 90.95 84.99 83.94 86.36 85.13 88.01
AnomalyTrans 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 84.14 96.88 90.06 89.79 100.0 94.62 94.48
DCdetector 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 83.04 89.22 86.02 93.25 100.0 96.51 94.26
PatchAD 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 83.06 90.51 86.62 93.28 100.0 96.52 94.60
Table 9: Parameter studies on project head constraint results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
Contraint Acc P R F1 Acc P R F1 Acc P R F1
0.0 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 97.42 98.03 97.72
0.1 98.92 91.96 98.33 95.03 99.04 94.43 98.33 96.34 98.72 97.35 98.04 97.70
0.2 98.92 92.05 98.20 95.02 99.11 94.42 98.91 96.61 98.91 97.43 98.66 98.04
0.3 98.89 92.03 97.93 94.88 99.14 94.41 99.11 96.71 98.89 97.42 98.59 98.00
0.4 98.86 92.09 97.54 94.73 99.12 94.45 98.97 96.66 98.71 97.42 97.94 97.68
0.5 98.96 92.12 98.60 95.25 99.02 94.40 98.17 96.25 98.73 97.42 98.03 97.72
0.6 98.92 92.11 98.20 95.06 99.01 94.39 98.07 96.19 98.92 97.45 98.70 98.07
0.7 98.84 91.91 97.59 94.67 99.07 94.43 98.55 96.44 98.76 97.39 98.14 97.76
0.8 98.91 92.01 98.20 95.00 99.10 94.40 98.81 96.55 98.92 97.42 98.73 98.07
0.9 98.78 91.99 96.90 94.38 98.96 94.37 97.70 96.01 98.92 97.47 98.64 98.05
Table 10: Parameter studies on anomaly threshold $\sigma$ results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
AR Acc P R F1 Acc P R F1 Acc P R F1
0.5 97.93 95.39 84.39 89.55 99.15 96.46 96.91 96.68 98.45 98.61 95.77 97.17
0.6 98.27 94.70 88.54 91.52 99.15 95.84 97.58 96.70 98.77 98.33 97.20 97.76
0.7 98.17 93.76 88.54 91.07 99.09 95.22 97.77 96.48 98.70 98.07 97.20 97.64
0.8 98.45 93.19 92.04 92.61 99.15 94.73 98.86 96.75 98.95 97.94 98.29 98.11
0.9 98.45 92.36 92.93 92.64 99.01 94.07 98.49 96.23 98.92 97.68 98.45 98.06
1.0 98.84 91.80 97.73 94.67 99.01 94.05 98.49 96.22 98.70 97.30 98.03 97.66
1.1 98.73 91.22 97.36 94.19 98.85 92.76 98.69 95.63 98.68 97.06 98.20 97.63
1.2 98.64 90.45 97.36 93.78 98.76 92.19 98.69 95.33 98.62 96.85 98.20 97.52
Table 11: Parameter studies on $d_{model}$ size results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
$d_{model}$ Acc P R F1 Acc P R F1 Acc P R F1
10 98.62 91.81 95.42 93.58 99.04 94.42 98.28 96.31 98.87 97.38 98.56 97.97
20 98.92 92.08 98.20 95.04 98.98 94.33 97.88 96.07 98.73 97.37 98.05 97.71
30 98.89 92.06 97.94 94.91 99.10 94.43 98.80 96.57 98.87 97.43 98.52 97.97
40 98.91 92.03 98.20 95.02 99.13 94.49 98.95 96.67 98.73 97.42 98.03 97.72
50 98.93 92.15 98.20 95.08 99.11 94.45 98.89 96.62 98.90 97.44 98.63 98.03
60 98.91 92.03 98.20 95.02 99.13 94.44 98.98 96.66 98.94 97.48 98.74 98.10
70 98.87 92.08 97.67 94.79 99.05 94.40 98.43 96.37 98.97 97.48 98.83 98.15
80 98.77 91.99 96.70 94.29 99.03 94.38 98.24 96.27 98.89 97.44 98.58 98.00
90 98.92 92.06 98.20 95.03 99.05 94.38 98.42 96.35 98.89 97.46 98.56 98.01
Table 12: Parameter studies on encoder layers $L$ results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
Encoder layer Acc P R F1 Acc P R F1 Acc P R F1
1 98.60 91.72 95.27 93.46 98.76 94.29 96.11 95.19 98.10 97.28 95.82 96.54
2 98.88 91.96 97.93 94.85 99.01 94.42 98.09 96.22 98.79 97.44 98.20 97.82
3 98.92 92.07 98.20 95.03 99.10 94.41 98.79 96.55 98.84 97.47 98.36 97.91
4 98.88 92.19 97.67 94.85 99.11 94.42 98.91 96.61 98.95 97.41 98.83 98.12
5 98.83 91.95 97.41 94.60 99.13 94.46 99.02 96.69 98.87 97.44 98.52 97.98
Table 13: Parameter studies on window size results (window size=105). All results are in %.
Dataset MSL SMAP PSM
Window size Acc P R F1 Acc P R F1 Acc P R F1
30 98.11 91.47 90.52 90.99 98.91 94.38 97.23 95.78 98.88 97.48 98.51 97.99
45 97.98 91.23 89.44 90.33 98.93 95.55 96.14 95.84 98.67 97.30 97.92 97.61
60 98.33 92.05 92.16 92.10 98.86 94.18 97.09 95.62 98.75 97.59 97.93 97.76
75 98.51 91.79 94.31 93.03 98.93 94.27 97.60 95.90 98.72 97.35 98.06 97.70
90 98.71 91.92 96.24 94.03 99.06 94.34 98.55 96.40 98.92 97.44 98.71 98.07
105 98.92 92.05 98.20 95.02 99.08 94.44 98.60 96.48 98.66 97.38 97.79 97.59
120 98.86 92.10 97.55 94.75 99.13 94.47 98.98 96.67 98.79 97.53 98.13 97.83
135 98.65 91.85 95.67 93.72 99.07 94.46 98.50 96.44 98.88 97.38 98.60 97.99
150 98.71 91.83 96.30 94.02 99.12 94.48 98.91 96.64 98.66 97.43 97.72 97.58
165 98.61 91.85 95.27 93.53 99.09 94.50 98.65 96.53 98.73 97.44 97.98 97.71
180 98.53 91.73 94.55 93.12 99.01 94.35 98.18 96.23 98.70 97.55 97.78 97.66
195 98.77 92.02 96.70 94.31 98.96 94.35 97.74 96.01 98.87 97.44 98.50 97.97
210 98.81 91.97 97.17 94.50 98.83 94.30 96.71 95.49 98.92 97.45 98.69 98.06
Table 14: Parameter studies on multi-scale patch size results (window size=105). All results are in %.
Dataset MSL SMAP PSM
Patch size Acc P R F1 Acc P R F1 Acc P R F1
[1] 98.49 91.73 94.12 92.91 99.09 94.47 98.69 96.53 98.66 97.21 97.99 97.60
[3] 98.57 91.67 95.08 93.34 99.08 94.44 98.61 96.48 98.84 97.56 98.26 97.91
[5] 98.34 91.70 92.66 92.18 99.02 94.40 98.20 96.26 98.83 97.44 98.37 97.91
[7] 98.45 91.96 93.41 92.68 99.13 94.46 99.04 96.70 98.88 97.43 98.56 97.99
[1, 3] 98.63 91.87 95.42 93.61 99.11 94.48 98.86 96.62 98.78 97.46 98.16 97.81
[1, 5] 98.63 91.88 95.42 93.61 99.07 94.43 98.58 96.46 98.72 97.37 98.04 97.70
[1, 7] 98.63 91.89 95.42 93.62 99.04 94.43 98.32 96.33 98.71 97.39 97.99 97.69
[3, 5] 98.78 91.99 96.90 94.38 99.07 94.39 98.63 96.46 98.91 97.43 98.66 98.04
[3, 7] 98.68 91.92 95.93 93.88 99.09 94.40 98.76 96.53 98.87 97.41 98.55 97.97
[5, 7] 98.61 91.86 95.27 93.53 99.08 94.44 98.63 96.49 98.85 97.40 98.46 97.93
[1, 3, 5] 98.85 92.01 97.54 94.69 99.12 94.47 98.92 96.64 98.82 97.37 98.39 97.88
[1, 3, 7] 98.74 92.00 96.46 94.18 99.09 94.40 98.71 96.50 98.79 97.43 98.21 97.82
[1, 5, 7] 98.92 92.05 98.20 95.02 99.04 94.38 98.36 96.33 98.67 97.36 97.87 97.61
[3, 5, 7] 98.56 91.76 94.89 93.30 99.08 94.42 98.61 96.46 98.87 97.45 98.52 97.98
[1, 3, 5, 7] 98.62 91.78 95.42 93.56 99.08 94.43 98.61 96.48 98.92 97.43 98.71 98.06
Appendix H Limitations
Although PatchAD achieves optimal results on multiple datasets, some limitations still need to be addressed. In order to ensure the optimization of different metrics, subtle adjustments to potential parameters are necessary. For instance, selecting an appropriate anomaly threshold is crucial to balance precision and recall. Developing an automated threshold adjustment scheme is a potential avenue for future research. Time series anomaly detection employs a wide range of evaluation metrics, each with its own emphasis, leading to potential biases in evaluating different algorithms. In the future, we aim to propose a unified set of evaluation metrics that consider multiple factors, contributing to the advancement of the community.
Appendix I Broader Impacts
Time series anomaly detection has significant academic implications and broader impacts in various domains. In finance, it aids in detecting fraudulent activities like credit card fraud and insider trading, safeguarding financial systems and preventing losses. In healthcare, it aids in early disease detection and monitoring, enabling timely interventions and improved patient outcomes. Overall, time series anomaly detection improves decision-making, risk mitigation, and system performance across multiple domains.