### v6

# PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection

Abstract
Time series anomaly detection is a pivotal task in data analysis, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios. While prior studies have largely employed reconstruction-based approaches, which limit the models’ representational capacities. Moreover, existing deep learning-based methods are not sufficiently lightweight. Addressing these issues, we present PatchAD, our novel, highly efficient multiscale patch-based MLP-Mixer architecture that utilizes contrastive learning for representation extraction and anomaly detection. With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 0.403M parameters. Its efficacy is demonstrated by state-of-the-art results across 8 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms. PatchAD significantly improves the classical F1 score by 6.84%, the Aff-F1 score by 4.27%, and the V-ROC by 2.49%. Simultaneously, an in-depth analysis of the mechanisms underlying PatchAD has been conducted from both theoretical and experimental perspectives, validating the design motivations of the model. The code is publicly available at https://github.com/EmorZz1G/PatchAD.
Index Terms:
Time series, anomaly detection, data mining, lightweight, MLP.
I Introduction
Time series anomaly detection (TSAD) is a pivotal aspect of data mining, focused on identifying abnormal patterns within time series data that significantly deviate from expected behaviour [1, 2, 3, 4]. In the era of big data, with the rapid advancement of large-scale sensing technologies and the continuous enhancement of storage capabilities, TSAD has been widely applied in diverse sectors. For instance, the Internet of Things (IoTs) serve to monitor abnormal events in sensor data [5, 6]. Unsupervised algorithms, capable of operating without labels, have gained wide popularity [7]. These include reconstruction-based [8, 9], density-based [10], contrastive learning [11, 12], autoregression-based [13, 14, 15], large language models (LLMs) [16] etc. However, extracting anomalies from extensive and complex temporal data presents significant challenges for different algorithms. Challenge 1: Accurately defining anomaly representations is challenging due to real-world anomalies’ diverse and dynamic nature. Effective anomaly detection requires models that can learn the inter-relationships between various data channels. Challenge 2: There is still significant potential for improvement in terms of training and inference speed, model complexity, and the lack of lightweight features. Challenge 3: In an unsupervised setting, the challenge lies in establishing a reliable basis to model normal and anomalous data. Specifically, several algorithms [17, 11, 18, 19] have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures and model degradation issue of contrastive learning. For instance, Xu et al. [17] devised a contrastive learning approach that leverages prior correlation and series correlation mining to elucidate the relationship between normal and anomalous patterns. However, this necessitated the employment of a maximization-minimization optimization strategy to train the model and prevent its degradation. Similarly, Yang et al. [11] extended the paradigm of image contrastive learning established by BYOL [20] by designing different temporal views to circumvent the challenges associated with negative sample construction in contrastive learning. Nonetheless, there exists the potential for performance degradation of the model when the patch settings are not appropriately configured.
Additionally, they either do not explicitly model inter-channel dependencies [11] or rely solely on basic MLPs or Transformers that focus on temporal information [18], mainly considering point-level features [19]. However, with the advancement of TSAD evaluation metrics, an increasing number of interval-based evaluation indicators [21, 22, 23, 24] have garnered research attention. Nevertheless, prior studies predominantly concentrated on the F1 with point ajustment [17], thereby neglecting the detection of interval anomalies. Consequently, there remains considerable potential for enhancing the detection capabilities for interval anomalies in current research [24].
Furthermore, the previous TSAD algorithms based on contrastive learning [17, 11, 12, 18] lack in-depth exploration from a theoretical perspective and fail to provide more valuable intuitions for future research. In contrast, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points. This difference aids our model in detecting anomalies, ensuring that normal data maintains consistent representations across different views, whereas anomalies do not. This fundamental principle forms the basis of our proposed anomaly detection method. Simultaneously, to further explore this concept, we conducted both experimental and theoretical analyses to investigate the underlying principles of this approach, thereby addressing the challenges from various perspectives.
In response to the above challenges and principles, we introduce PatchAD, a lightweight Patch-based MLP architecture for time series Anomaly Detection. Specifically, proposed Multi-scale Patching and Embedding decomposes the input data along the temporal dimension into non-overlapping patches, extracting semantically richer features beyond simple point-level characteristics. Besides, various submodels address various time scales, each focusing on different patch sizes. Patch Mixer Encoders are introduced to learn the complex relationships within and between patches and channels. A shared MixRep Mixer uniformly represents the feature space, while a Dual Project Constraint prevents the model from adopting overly simplistic solutions, thereby enhancing its representational capabilities. At the same time, the lightweight model configuration enables PatchAD to better tackle Challenge 2, while the intricate encoder design enhances the model’s feature learning capacity to confront Challenge 1.
In detail, this work makes the following contributions:
Figure 1: Overall performances under different metrics.
Figure 2: Model Params, FLOPs, and Latency.
(1) State-of-the-art Anomaly Detection (Figure2): We propose a novel contrastive learning-based time series anomaly detection method. PatchAD demonstrates exceptional performance across eight benchmark datasets, exceeding 30 baselines across a range of evaluation metrics, including commonly used metrics and interval anomaly indicators.
(2) Lightweight and Efficient Design (Figure2): PatchAD employs a multi-scale patch-based MLP-Mixer architecture with different novel strategies, achieving remarkable efficiency and compactness compared to Transformer-based models. Moreover, we conducted an analysis of the model’s scalability from both experimental and theoretical perspectives.
(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by practical and theoretical analyses that elucidate the model’s underlying principles and design choices.
II Preliminaries
II-A Problem Formulation
In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting: $\mathcal{X}=(x_{1},x_{2},\cdots,x_{T})$, $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines. The objective is to train a model using $\mathcal{X}_{train}$ and provide predictions $\mathcal{Y}_{test}=(y_{1},y_{2},\cdots,y_{T^{\prime}})$ on another sequence $\mathcal{X}_{test}$ of length $T^{\prime}$. Here, $y_{t}\in\{0,1\}$, where 1 signifies an anomaly at that time point and 0 indicates a normal point.
II-B Time Series Anomaly Detection
Figure 3: The workflow of the proposed PatchAD framework.
The time series anomaly detection methods encompass statistical approaches, classical machine learning methods, and deep learning techniques [25]. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) [26] and Support Vector Data Description (SVDD) [27]. However, machine learning methods often struggle to extract meaningful features and detect anomalies when confronted with high-dimensional, complex, and voluminous time series data, ultimately resulting in suboptimal performance [4, 28].
Deep learning methods include OmniAnomaly [9], USAD [19], LSTM-VAE [8], and DeepAnT [29], which are trained on normal data and detect anomalies by inference using reconstruction errors. They often rely on reconstruction assumptions, wherein the model’s ability to restore anomalous data diminishes after being trained on normal data. Previous studies [30, 31] have indicated that this prior becomes less applicable when the model’s reconstruction capability is excessively strong, leading to a decline in anomaly detection performance.
Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [32]. While [33] designed six different anomaly injection methods to train the model in a self-supervised approach, [34] utilized outlier exposure and mixup to construct a self-supervised task. However, due to the distributional discrepancies between real and synthetic data, this results in subpar performance when confronted with real-world data, as observed in [33, 34].
Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like [35] introduced sequence variational autoencoder model based on smoothness constraints, whereas [17] proposed prior and sequence associations for temporal modeling, both achieving enhanced performance. Recently approaches based on diffusion models and Large language model have also been proposed [36, 16]. Although significant advancements have been made in the performance of time series anomaly detection methods, most of these approaches are based on Transformers or sophisticated attention models [17, 11, 36, 16]. This reliance results in models that are not sufficiently lightweight and exhibit slower computational efficiency, rendering them unsuitable for certain real-world scenarios that require rapid detection.
II-C MLP Mixer & Contrastive Learning for Time Series Analysis
The MLP Mixer initially proposed as a novel architecture for computer vision [37], has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like [38, 39, 40, 41] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. For example, PatchTST [39] employs a single-scale patch-based MLP as its core architecture, whereas the work in [41] applies an MLP to capture frequency-domain information in time series data for predictive modeling. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time series forecasting (LSTF). Constrained by space, further detailed descriptions about Mixer are provided in Appendix -A1.
Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs [20, 42, 12]. Additionally, methods like BYOL [20] and SimSiam [42] have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as [12] and [43] have explored integrating contrastive learning into time series representation learning. TS2Vec [44] performs contrastive learning in a hierarchical way over augmentation views. [17, 11], on the other hand, combine self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection. In consideration of space limits, technical details are provided in Appendix -A2.
III Proposed Method
III-A Overall Architecture
Figure 3 illustrates the overall structure of PatchAD, which adopts patch sizes at multiple scales. We exemplify its architecture using a single scale. PatchAD is an $L$-layer network, where each layer comprises distinct modules: MLP Mixers, Project head, and ReWeight, forming its structure. We have devised four distinct types of MLP Mixers: Channel Mixer, Inter Mixer, Intra Mixer, and MixRep Mixer, each dedicated to learning various channels, inter-patch, intra-patch relationships, and a unified representation space. The Project head serves to prevent the model from converging to trivial solutions, while the ReWeight module manages the weights across different layers.
The focal point of our design lies in the four distinct MLP Mixers. We utilize various MLP Mixers to learn different representations from the input. Normal points share a common latent space across different views, while anomalies, being scarce and lacking specific patterns, struggle to share a coherent representation with normal points. Consequently, normal points exhibit minor discrepancies across different views, whereas anomalies showcase larger variations. PatchAD leverages inter-intra discrepancy to model this relationship.
Our anomaly detection approach, PatchAD, leverages the principle that normal time series points, characterized by low information entropy, exhibit consistent representations across diverse perspectives. In contrast, anomalies with higher entropy struggle to maintain such consistency (Figure 3(c)). PatchAD employs an Inter-Patch Encoder to capture detailed features and an Intra-Patch Encoder to learn generalized trends. This combined approach enables a comprehensive understanding of normal and anomalous representations, achieving effective anomaly detection.
III-B Multi-Scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. Patching divides the time series window into smaller patches [39]. Multi-scale Patching enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$ [39]. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ (\mathbb{R}^{C\times T}\rightarrow\mathbb{R}^{C\times N\times P}),$ (1)
where $\operatorname{PE}(\cdot)$ is positional embedding [17]. PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. To make feature extraction and learning easier, these are used to embed the intra-patching and inter-patching information into the $D$-dimensional space, respectively. And $\operatorname{VE}(\cdot)$ is a single linear layer.
$\mathcal{N}$ $=\operatorname{VE}(\mathcal{X}),\ (\mathbb{R}^{C\times N\times P}\rightarrow\mathbb{R}^{C\times N\times D}),$ (2)
$\mathcal{P}$ $=\operatorname{VE}(\mathcal{X}),\ (\mathbb{R}^{C\times N\times P}\rightarrow\mathbb{R}^{C\times P\times D}).$
For the sake of brevity and clarity, the above describes the process of single-scale patching. In the last part of this chapter, we describe how multi-scale patching is implemented (see Sec. III-G).
III-C Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. The Multi-Layer Perceptron (MLP) consists of two fully connected (FC) layers, a normalization layer and a ReLU activation, and utilizes a residual connection from input to output. The MLP can be represented as:
$\operatorname{MLP}(\mathcal{X})=\operatorname{FC}(\operatorname{ReLU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (3)
The MLP Mixers of different types utilize MLP to model the feature variations along different dimensions, which can be represented as:
$\operatorname{Mixer}(\mathcal{X},dim=i)$ (4)
$=\operatorname{Transpose}(\operatorname{MLP}(\operatorname{Transpose}(\mathcal{X},\text{dim = i})),\text{dim = i}),$
where $\operatorname{Transpose}(\cdot,\text{dim = i})$ denotes the operation of swapping the $i$-th dimension with the last dimension. The purpose of this is to enable the MLP to effectively capture the features along different dimensions. The following are the different types of MLP Mixers utilized in PatchAD:
  1. 1.
    Channel Mixer captures correlations among different channels.
  2. 2.
    Inter Mixer captures inter-patch representations, improving detailed expressions.
  3. 3.
    Intra Mixer captures intra-patch representations, improving generalized expressions.
  4. 4.
    MixRep Mixer embeds two distinct views into the same representation space.
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation. Specifically, for each layer of the Patch Mixer Encoder, there exist two types of input features: inter-patch, denoted as $\mathcal{N}$, and intra-patch, represented as $\mathcal{P}$.
Next, we elaborate on the workflow of four types of mixers in detail. Initially, these features are processed by the shared-parameter Channel Mixer, expressed as follows:
[…]
Similarly, the Inter Mixer and Intra Mixer utilize the outputs of the channel mixer as their respective inputs, yielding:
[…]
Here, $\operatorname{InterMixer}$ and $\operatorname{InraMixer}$ denote the functions of the Inter Mixer and Intra Mixer, respectively. It is important to note that although both operate on feature variation at dim=2, the differing shapes of the input tensors result in $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$. Consequently, $\operatorname{InterMixer}$ processes $\mathcal{N}$ along dimension $N$ to learn detailed features of patches, whereas $\operatorname{IntraMixer}$ processes $\mathcal{P}$ along dimension $P$ to acquire global knowledge of patches. Secs. V-A and V-B provide a detailed experimental and theoretical analysis of the nature of learning in both cases.
Subsequently, to enable the model to learn low-dimensional invariant information from different types of features and thereby enhance the generalization capability of PatchAD, inspired by [45], the MixRep Mixer is employed to embed the two distinct types of features into a unified representation space, expressed as follows:
[…]
$\mathcal{N}=\operatorname{MixRepMixer}(\operatorname{InterMixer}(\operatorname{ChannelMixer}(\mathcal{N}))),$ (8)
$\mathcal{P}=\operatorname{MixRepMixer}(\operatorname{IntraMixer}(\operatorname{ChannelMixer}(\mathcal{N}))).$
In summary, the four mixer workflows correspond to Figure 3(b).
In the end, two separate simple MLPs, e.i., $\operatorname{Rec_{1}}$ and $\operatorname{Rec_{2}}$ are utilized to reconstruct the original data, which are then combined.
$\hat{\mathcal{X}}_{1}=\operatorname{Rec_{1}}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec_{2}}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (9)
Simultaneously, the structures of $\operatorname{Rec_{1}}$ and $\operatorname{Rec_{2}}$ are identical. Initially, the last two channels of $\mathcal{N}$ and $\mathcal{P}$ are flattened using $\operatorname{Flatten}(\cdot)$. Subsequently, the features are reconstructed into the original time series $\mathcal{X}$ through a MLP. This process can be represented as:
$\operatorname{Rec_{1}}(\mathcal{N})$ $=\operatorname{MLP}(\operatorname{Flatten}(\mathcal{N})),$ (10)
$\operatorname{Rec_{2}}(\mathcal{P})$ $=\operatorname{MLP}(\operatorname{Flatten}(\mathcal{P})).$
III-D Dual Project Constraint
To prevent the model from converging to trivial solutions, each layer’s output requires to be constrained by Dual Project Head. Prior research suggests that employing a similar structure effectively prevents models from converging to trivial solutions. For instance, SimCLR [45] uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training. The projection head comprises two layers of MLP connections, excluding non-linear and normalization layers. Hence, for both the inter-patch and intra-patch encoders, we can derive the projected representations $\mathcal{N}^{\prime}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}^{\prime}\in\mathbb{R}^{C\times P\times D}$. This computation can be represented as:
$\mathcal{N}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{N})}),\ \mathcal{P}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{P})}).$ (11)
To amalgamate outputs from different layers and prevent model degradation, we employ a simple ReWeight module to assign distinct weights to various layers. Considering representations from $L$ layers, $\{\mathcal{N}_{1},\mathcal{N}_{2},\cdots,\mathcal{N}_{L}\}$ and $\{\mathcal{P}_{1},\mathcal{P}_{2},\cdots,\mathcal{P}_{L}\}$, we can derive the final result:
$\mathcal{N}_{l}$ $=\alpha_{l}\cdot\mathcal{N}_{l},\ \alpha_{l}=\operatorname{Softmax}(\alpha),$ (12)
$\mathcal{P}_{l}$ $=\beta_{l}\cdot\mathcal{P}_{l},\ \beta_{l}=\operatorname{Softmax}(\beta).$
Therefore, the outputs of the encoder can be represented as:
$\mathcal{N}=\alpha_{1}\cdot\mathcal{N}_{1}+\cdots+\alpha_{L}\cdot\mathcal{N}_{L},$ (13)
$\mathcal{P}=\alpha_{1}\cdot\mathcal{P}_{1}+\cdots+\alpha_{L}\cdot\mathcal{P}_{L}.$
The introduction of ReWeight allows for a better balance between outputs from different layers, resulting in improved performance.
III-E Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both. In Sec. IV-C1, the reasons for choosing the KL divergence as the basic loss are analyzed. The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$ (14)
$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
where $\operatorname{KL}(\cdot||\cdot)$ represents the KL divergence distance, and $\operatorname{StopGrad}$ denotes gradient stoppage, employed for asynchronously optimizing the two branches. The combined loss is defined as:
$\mathcal{L}_{cont}=\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}.$ (15)
Due to the mismatch in dimensions between $\mathcal{N}$ and $\mathcal{P}$, an initial upsampling of both is required for comparability. For the inter-patch view, where only differences between patches exist, we replicate the final $\mathcal{N}$ within patches. Conversely, for the intra-patch view, exhibiting anomalies between points within patches, we replicate it multiple times to obtain the final $\mathcal{P}$. This also addresses the necessity for introducing multi-scale, as it compensates for the information loss during upsampling.
To prevent model degradation, we augment the original loss function with an additional constraint from the projection head, defined as:
$\mathcal{L}_{proj}=\frac{\mathcal{L}_{\mathcal{N}^{\prime}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}+\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}^{\prime}}}{len(\mathcal{N})}.$ (16)
The final PatchAD loss comprises the two aforementioned components and the reconstruction term ($\mathcal{L}_{rec}=\operatorname{MSE}(\hat{\mathcal{X}},\mathcal{X})$), which is accompanied by a constraint coefficient to regulate the strength of the constraint.
$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}+\mathcal{L}_{rec}.$ (17)
III-F Anomaly Score
Based on the assumption that normal points across different views share hidden patterns, while anomalies exhibit larger discrepancies, the difference between the two views can be employed as an anomaly score. The final anomaly score can be expressed as:
$\operatorname{AnomalyScore}(\mathcal{X})=\sum\operatorname{KL}(\mathcal{N},\mathcal{P})+\operatorname{KL}(\mathcal{P},\mathcal{N}).$ (18)
Based on the aforementioned anomaly score, we set a hyperparameter $\sigma$ to determine whether a point is considered an anomaly. If it surpasses the threshold, it is identified as an outlier.
III-G Multi-Scale of PatchAD
For single-scale PatchAD, it features a comprehensive Patching, Patch Mixer Encoder, and Dual Project. This means that, assuming the patch size for single-scale PatchAD is $P_{i}$, we obtain the final output of the Patch Mixer Encoder, $\mathcal{\hat{X}}_{i}$, as well as the final outputs of the Dual Project, $\mathcal{N}_{i}$ and $\mathcal{P}_{i}$. These outputs are used to update the loss function $\mathcal{L}_{i}$ (as shown in Eq. (17)).
Here, $i$ denotes the $i$-th single scale. For a multi-scale PatchAD, there are typically $M$ branches. We optimize the loss functions $\{\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{M}\}$ in an integrated manner, allowing different branches to learn features at various scales. During the testing phase, each single-scale branch produces its own anomaly score $AS_{i}$, and the final anomaly score is the average of the scores from all branches, given by $AS=\frac{1}{M}\sum_{i=1}^{M}AS_{i}$.
IV Experiments
IV-A Setups and Evaluations
IV-A1 Datasets
We evaluated different algorithms using a total of 8 datasets, with details referred to in the appendix.
IV-A2 Metrics
Considering comprehensively, we adopt two groups of metrics to conduct extensive experiments. Detailed explanations for why these metrics are used can be found in Appendix -D3.
  1. 1.
    Metrics using point adjustment (PA): We prioritize these metrics because they are now the most commonly utilized group. To compare with advanced algorithms (including but not limited to [11, 17, 9, 46, 19, 15]) that used point-adjusted metrics for evaluation, we employ Acc, Pre, Rec, and PA-F1 as the main metrics.
  2. 2.
    Metrics without point adjustment: Meanwhile, as many current studies [28, 24] are gradually using classic metrics (AUC, F1-cls) and interval evaluation metrics (Aff-F1, VUS), in order to compare with these algorithms (including but not limited to [33, 47, 16, 48, 49, 50, 11]), we organize additional experiments for further analysis in Sec. IV-B2.
IV-A3 Baselines
We compared our model against 30+ benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) The classic methods; (6) Change point detection and time series segmentation methods; (7) Contrastive learning; (8) Large Language Model; (9) Diffusion model.
The details regarding the selection of the algorithms are further discussed in Appendix -D4.
IV-B Comparison Results
IV-B1 Comparisons via Metrics with PA
From Table I, PatchAD achieves the best average-level PA-F1 performance. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive. In the evaluation of the PA-F1 metric, DCdetector and AnomTrans emerge as the state-of-the-art(SOTA) models, with their robust capability to capture local anomalies granting them a competitive edge over other models in PA-F1. In contrast, PatchAD enhances its applicability to local anomalies through adaptive learning of local feature representations, achieving SOTA performance. Compared with AnomTrans’s limited temporal granularity, PatchAD’s multi-scale approach can better capture anomalies of different lengths, thus achieving superior PA-F1 performance.
Simultaneously, we observe that the additional models do not perform exceptionally well on the PA-F1 metric, primarily due to two reasons: firstly, they are mainly designed for the detection of interval anomalies, a fact that becomes evident in the subsequent analysis. Secondly, the PA-F1 metric employs point adjustment techniques, which can result in some undetected anomalies being classified as true positives, thereby introducing a potential inflation effect in the metric [51]. While PA-F1 is suitable for comparing the performance of different models in research contexts, it is less appropriate for assessing model performance in real-world scenarios. Consequently, this also accounts for the challenge faced by PatchAD in distinguishing itself from previous models. Therefore, in the following section, a new set of metrics will be employed to analyze the models from different perspectives.
TABLE I: Comprehensive comparative results on four datasets. The P, R and PA-F1 are the precision, recall and F1 with PA. All results are in %, the best in Bold, and the second in underlined.
Dataset MSL SMAP PSM SWAT Average PA-F1
Metric P R PA-F1 P R PA-F1 P R PA-F1 P R PA-F1
LOF 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
OCSVM 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
U-Time 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
IForest 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
DAGMM 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
ITAD 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
VAR 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
Deep-SVDD 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
BOCPD 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
LSTM-VAE 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BeatGAN 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetetor 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
TimesNet 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
PatchTST 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
GPT2-Adapter 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
NPSR-pt [48] 10.60 100.0 19.16 24.99 96.37 39.68 27.74 100.0 43.43 12.29 100.0 21.89 31.04
NPSR-seq [48] 17.66 100.0 30.02 12.79 100.0 22.68 37.87 99.95 54.81 86.76 88.58 87.66 48.80
M2N2 [52] 89.38 92.48 90.90 87.37 72.15 79.04 70.51 89.99 79.07 99.42 70.96 82.82 82.96
AdaMemBLS [4] 88.01 93.20 90.53 88.01 93.20 90.53 89.59 95.45 92.43 93.19 82.78 87.68 90.29
PatchAD 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
(a) Global point
(b) Contextual point
(c) Seasonal
(d) Group point
(e) Trend
Figure 4: Visualization of generated synthetic data, PatchAD anomaly scores and ground truth.
IV-B2 Comparisons via Metrics without PA
Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (MSL, SMAP, SWAT, WADI, PSM, and SWAN). Here, Aff-F1 is the affiliation metric [22]. V-ROC and V-RR are volumes under the surfaces created based on ROC curve and PR curve [21], respectively. In terms of the F1-cls, Aff-F1, and V-ROC metrics, PatchAD demonstrates relative performance improvements of 6.84%, 4.27%, and 2.49% compared to the second-place model, respectively. This indicates a significant enhancement in PatchAD’s capacity for detecting interval anomalies compared to previous models. Notably, we observe that AnomTrans and DCdetector, which perform well on the PA-F1 metric, exhibit poor performance on interval anomaly metrics. This is primarily because these models are more focused on local or point anomalies, resulting in better performance on the PA-F1 metric. PatchAD, on the other hand, benefits from additional learning of generalized global features, enhancing its performance on interval anomalies, thus improving both PA-F1 and Aff-F1 metrics.
Furthermore, we note that although TimesNet was proposed earlier, it currently exhibits relatively strong performance when considering the two sets of metrics. This can be attributed to the diversity of contemporary anomalies, which vary in type, periodicity, and length. TimesNet’s capability to learn temporal information across different frequencies enables it to handle various anomalies more effectively. PatchAD similarly adopts this approach, with multi-scale temporal feature extraction ensuring the learning of information across different temporal frequencies.
NPSR also takes into account the utilization of two distinct feature types in its design and uniquely processes anomaly scores, thereby ensuring its excellent theoretical lower bound, which guarantees good performance in the face of real interval anomalies. Likewise, PatchAD benefits from the assurance provided by the lower bound of information entropy and the diverse features extracted through different Mixers, resulting in similarly outstanding performance. In contrast, COUTA’s training relies on anomaly injection, leading to varying adaptability across different real datasets; for instance, it performs well on the PSM and SWAN datasets but shows average performance on the WADI and SWAT datasets.
TABLE II: Comprehensive comparative results on six datasets. All results are in %, the best in Bold, and the second in underlined.
Datasets MSL SMAP SWAT
Methods AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR
LOF [2000] 55.75 19.36 66.86 60.85 17.19 62.39 23.64 58.74 57.56 18.59 84.59 58.22 5.25 66.67 41.88
IForest [2008] 59.09 10.90 70.44 64.74 18.29 61.15 23.78 50.52 59.31 16.45 83.80 38.52 5.27 72.29 52.07
PCA [2003] 53.25 10.33 70.72 59.42 18.08 58.70 22.96 50.72 43.90 11.69 81.85 26.05 5.25 61.87 44.75
Deep SVDD [2018] 61.80 24.30 68.98 59.54 19.36 61.14 21.28 59.64 41.95 11.56 86.81 22.57 69.77 50.96 56.48
USAD [2020] 57.12 22.50 67.69 58.86 19.77 62.79 16.57 50.41 33.78 11.28 88.66 21.70 69.28 84.38 62.74
TCN-ED [2021] 51.14 19.66 67.93 58.62 18.13 58.91 22.90 67.89 43.79 11.73 89.23 21.65 69.26 61.53 57.94
TranAD [2022] 50.03 22.81 68.97 57.25 17.55 59.74 22.74 50.51 45.15 11.98 88.90 25.50 71.42 86.25 64.19
NCAD [2022] 60.20 22.05 67.25 67.29 21.32 53.45 23.09 68.32 48.37 13.22 82.92 68.72 73.89 76.76 44.91
Deep IF [2023] 55.94 19.11 67.94 52.33 17.04 60.09 29.14 69.59 56.89 14.77 89.52 21.65 69.26 50.19 56.39
AnomTrans [2021] 52.61 18.39 66.81 53.03 13.21 52.18 16.06 71.44 52.61 16.55 80.80 23.44 66.77 21.49 8.29
TimesNet [2022] 57.18 21.24 67.64 64.63 20.67 53.73 24.37 66.19 48.83 12.91 73.24 21.66 71.55 31.10 10.29
GPT2-Ada. [2023] 52.03 13.72 68.81 53.57 17.38 55.48 24.12 69.48 57.31 15.78 52.30 22.30 68.41 50.01 10.29
DCdetector [2023] 50.31 11.62 67.85 52.00 14.68 58.50 26.56 71.37 60.38 16.87 52.78 23.24 68.56 50.84 14.68
D3R [2024] 63.00 23.98 69.79 69.02 20.93 54.56 22.71 67.92 48.67 13.23 79.95 45.89 68.96 84.38 57.27
COUTA [2024] 55.59 20.88 67.68 61.49 18.68 58.74 22.69 67.97 44.72 11.91 75.54 47.65 46.92 72.57 25.06
NPSR [2023] 61.16 23.72 68.42 66.44 19.72 38.74 22.68 67.95 41.91 11.42 90.18 76.88 75.86 84.93 65.42
M2N2 [2024] 59.15 21.76 67.88 65.40 21.55 54.05 22.92 67.95 48.74 14.36 67.79 38.50 73.44 61.14 26.16
AdaMemBLS [2025] 51.78 19.11 71.41 57.46 16.06 53.66 26.93 69.49 55.57 14.26 81.61 74.10 69.76 82.00 66.12
PatchAD 60.94 24.92 70.10 66.43 19.65 60.27 24.74 70.95 48.30 13.71 91.19 77.69 78.76 89.91 65.74
Datasets WADI PSM NIPS-TS-Swan
Methods AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR
LOF [2000] 65.44 10.58 34.36 66.67 41.88 81.28 23.91 62.32 80.42 64.55 77.89 21.37 19.62 89.42 88.56
IForest [2008] 74.84 30.88 67.90 72.29 52.07 75.12 38.52 56.77 75.19 55.50 85.12 46.18 27.81 91.27 90.98
PCA [2003] 49.47 34.30 50.01 61.87 44.75 72.06 43.20 78.74 73.97 54.55 72.88 49.18 27.35 90.42 90.23
Deep SVDD [2018] 59.47 7.34 52.77 50.96 56.48 78.09 44.46 70.06 53.97 66.92 75.32 55.99 69.66 80.89 91.74
USAD [2020] 53.61 10.86 70.90 84.38 62.74 64.53 47.90 70.88 64.02 69.18 66.76 52.63 61.22 91.40 94.50
TCN-ED [2021] 53.01 10.86 70.90 61.53 57.94 63.97 43.46 69.42 57.48 42.64 71.37 49.22 71.03 80.88 91.73
TranAD [2022] 52.60 11.78 69.54 86.25 64.19 65.22 43.20 73.30 72.42 56.02 70.00 51.93 70.49 88.01 92.59
NCAD [2022] 62.84 10.87 70.96 76.76 44.91 61.76 46.61 63.10 67.09 48.26 52.24 37.57 68.87 85.31 93.44
Deep IF [2023] 51.43 10.86 70.90 50.19 56.39 69.06 43.47 69.41 67.71 45.98 73.30 49.19 71.05 80.89 91.74
AnomTrans [2021] 52.98 11.24 66.77 21.49 8.29 52.18 39.81 68.13 49.12 31.56 54.62 44.86 69.11 83.56 81.64
TimesNet [2022] 65.08 17.65 49.05 31.10 10.29 58.74 43.60 72.15 66.37 48.05 53.62 43.16 69.53 91.58 94.30
GPT2-Ada. [2023] 51.21 10.68 71.04 50.01 10.29 51.23 35.24 69.17 52.45 32.79 58.77 45.23 65.09 84.31 84.23
DCdetector [2023] 50.12 11.33 73.23 50.84 14.68 50.38 22.72 68.08 51.73 32.44 50.35 48.83 70.99 85.84 83.11
D3R [2024] 51.39 12.86 72.17 84.38 57.27 60.64 44.24 69.99 66.92 49.38 62.54 49.45 72.10 89.27 88.29
COUTA [2024] 50.52 26.92 37.44 72.57 25.06 69.36 48.21 73.21 64.50 43.15 71.38 47.19 66.54 82.15 92.26
NPSR [2023] 80.19 50.67 75.74 84.93 65.42 70.67 51.02 70.56 70.55 52.12 62.61 49.30 71.07 90.90 90.35
M2N2 [2024] 55.68 13.55 70.90 43.99 7.98 54.91 47.71 69.84 40.06 30.61 70.27 57.47 71.05 89.77 90.18
AdaMemBLS [2025] 53.85 29.09 70.93 63.47 21.40 63.26 46.69 71.51 65.07 49.24 51.27 49.35 71.09 87.12 85.92
PatchAD 85.27 52.82 79.04 88.29 49.60 64.74 49.33 76.73 65.11 40.72 79.53 64.52 72.36 95.62 91.52
IV-C Model Analysis
IV-C1 Ablation Studies
Ablation on Components: To study the impact of different components in our model, we conducted ablation experiments across four datasets, as presented in Table III. In a manner analogous to the comparative experimental setup, we selected two sets of metrics to analyze the impact of components on the model from different perspectives. The first set comprises point-adjusted metrics, including Acc, Pre, Rec, and PA-F1. These metrics exhibit superior adaptability to noisy labels and demonstrate greater robustness in the presence of mislabeled data [28, 24]. The second set consists of metrics that do not employ point adjustments, including AUC, F1-cls, V-ROC, and V-PR. While AUC and F1-cls exhibit poorer robustness to label noise and lower flexibility, their stricter evaluation criteria ensure a lower bound on the accuracy of the assessment [24]. Conversely, V-ROC and V-PR are two interval-based evaluation metrics that provide more accurate assessments for longer intervals of anomalies, thereby enhancing comprehensiveness. The “wo_ch_sharing” method employs two independent Channel Mixer modules to model channel features, while “wo_ch_mixer” denotes removing the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modelling inter-channel information and sharing channel information benefit model learning. Besides, “wo_pos_emb” represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, “wo_constraint” showcases the impact of removing the constraint on the projection head. In conjunction with Figure 5, we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
Considering the limitation of insufficient real labels for threshold selection in practical scenarios, we validated the practicality of PatchAD using the dynamic threshold algorithm based on Extreme Value Theory (SPOT) [53], as shown in the second-to-last row of Table III. Results indicate that its performance gap with the optimal model is minimal, demonstrating that dynamic threshold settings can replace manual $\sigma$ tuning to achieve favorable outcomes in real-world applications. It is important to note that AUC, V-ROC, and V-PR are threshold-agnostic metrics, so using SPOT does not affect these results.
Figure 5: Loss analysis of PatchAD.
TABLE III: Ablation studies on MSL, SMAP, WADI, and PSM datasets.
Dataset MSL SMAP WADI PSM
With PA Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1
wo_ch_sharing 98.78 91.93 96.9 94.35 98.84 93.5 97.75 95.58 98.71 84.99 94.25 89.38 97.69 97.62 93.96 95.75
wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.5 97.1 95.27 98.92 85.4 97.93 91.24 96.77 98.23 89.98 93.92
wo_mixrep_mixer 96.41 89.54 74.66 81.43 95.25 91.52 68.55 78.7 95.76 67.55 51.99 58.76 97.35 97.87 92.44 95.08
wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74 96.89 97.84 90.77 94.17
wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53 93.95 97.22 80.45 88.04
PatchAD with SPOT 98.49 89.94 96.46 93.08 98.93 94.2 97.65 95.89 98.53 84.18 91.53 87.7 97.75 97.79 93.99 95.85
PatchAD (Ours) 98.92 92.05 98.2 95.02 99.1 94.49 99.13 96.75 99.06 85.92 100 92.43 98.89 97.72 98.52 98.11
Dataset MSL SMAP WADI PSM
Without PA AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR
wo_ch_sharing 53.54 20.57 60.07 18.14 59.46 22.71 43.36 11.60 85.24 50.33 88.13 49.22 64.65 48.00 65.93 40.13
wo_ch_mixer 53.09 20.16 59.69 17.99 59.47 22.68 43.06 11.59 82.41 46.50 86.88 48.00 65.02 48.08 66.59 40.65
wo_mixrep_mixer 53.50 20.70 60.02 18.13 59.35 22.69 43.41 11.61 81.14 50.61 87.96 47.09 64.47 48.11 65.91 40.09
wo_pos_emb 52.91 19.55 59.68 17.95 55.95 22.68 46.87 12.15 84.63 51.02 87.36 49.96 64.76 48.04 66.17 40.25
wo_constraint 53.53 20.79 59.97 18.11 59.93 22.71 42.85 11.52 83.56 51.29 86.78 46.09 64.93 47.95 66.54 40.53
PatchAD with SPOT 60.94 23.45 66.43 19.65 60.27 23.77 48.30 13.71 85.27 52.75 88.29 49.60 64.74 48.58 65.11 40.72
PatchAD (Ours) 60.94 23.92 66.43 19.65 60.27 24.74 48.30 13.71 85.27 52.82 88.29 49.60 64.74 49.33 65.11 40.72
Ablation on Loss Functions: To facilitate the spatial separation of the two distinct views, the design in Eq. (14) employs the Kullback-Leibler ($\operatorname{KL}$) divergence as the foundational loss. To further analyze the impact of various types of loss functions on the final model performance, we selected the following three common and widely utilized distance or distributional metric functions as alternatives, enabling an examination of their effects on the ultimate outcomes.
  1. 1.
    L2 Distance (L2).
  2. 2.
    Wasserstein Distance (Wasserstein).
  3. 3.
    Jensen–Shannon Divergence (JSD).
Table IV illustrates the performance of the PatchAD model after substituting the three loss functions in comparison to the model utilizing KL divergence (Ours).
In the comparison of the first set of metrics employing PA, the models based on L2 and JSD exhibit average performance that falls short of the final model; however, the overall difference from Ours is minimal. Conversely, the Wasserstein distance yields a PA-F1 score of only 83.93 on the WADI dataset, indicating a performance that is significantly inferior to that of the other distance functions. This may be attributed to the fact that the Wasserstein distance contributes to anomaly scores in a more smooth manner, which, in turn, results in a reduced recall rate for the model and consequently leads to suboptimal performance.
In the comparison of the second set of metrics, PatchAD consistently ranks as the best across nearly all datasets and metrics. Additionally, models trained with the other three foundational distances exhibit markedly lower performance in AUC, V-ROC, and V-PR compared to PatchAD. This suggests that these models do not possess sufficiently clear decision boundaries for anomalies, thereby complicating the model’s ability to accurately differentiate between anomalies and normal instances.
Overall, we posit that JSD represents the best alternative to KL divergence, primarily because the construction of JSD is based on KL divergence, thereby retaining KL divergence’s property of optimizing over probability distributions. However, JSD is a symmetric loss function, while KL divergence is asymmetric; this discrepancy limits the ability to achieve asynchronous optimization of the parameters in different modules (the Inter Mixer and Intra Mixer). Ultimately, this results in JSD performing poorly in enhancing the distinction between normal and abnormal instances, a phenomenon that aligns with previous work [17].
TABLE IV: The Influence of different loss functions as fundamental distances on PatchAD.
Dataset MSL SMAP WADI PSM
With PA Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1
L2 98.12 88.92 93.82 91.30 98.90 93.75 97.94 95.80 99.03 86.89 97.93 92.08 97.12 97.65 91.84 94.65
Wasserstein 98.16 88.85 94.41 91.55 99.20 95.33 98.57 96.92 98.07 80.35 87.85 83.93 97.59 97.45 93.75 95.56
JSD 98.32 90.79 93.51 92.13 99.16 97.81 95.55 96.66 99.01 85.37 98.57 91.50 98.41 97.88 96.33 97.10
PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.49 99.13 96.75 99.06 85.92 100.00 92.43 98.89 97.72 98.52 98.11
Dataset MSL SMAP WADI PSM
Without PA AUC F1-cls V-ROC V-PR AUC F1-cls V-ROC V-PR AUC F1-cls V-ROC V-PR AUC F1 V-ROC V-PR
L2 53.52 20.10 60.24 18.15 58.69 22.68 44.01 11.67 85.54 50.22 87.44 47.71 64.64 48.17 65.17 40.12
Wasserstein 53.39 19.62 60.13 18.08 59.56 22.69 43.24 11.57 84.15 51.61 86.58 46.13 64.66 48.20 65.71 40.56
JSD 53.64 19.57 60.34 18.13 58.78 22.68 44.02 11.67 85.08 52.62 87.92 49.24 65.66 49.11 65.69 40.36
PatchAD (Ours) 60.94 23.92 66.43 19.65 60.27 24.74 48.30 13.71 85.27 52.82 88.29 49.60 64.74 49.33 65.11 40.72
IV-C2 Visual Analysis
We visualized PatchAD’s response to various anomalies in Figure 4. Synthetic univariate time series were generated by [54], incorporating global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies. The first row depicts the original data, while the second row displays PatchAD’s output of anomaly scores against the anomaly threshold. In particular, PatchAD has the ability to detect various types of anomalies. Figure 6 presents the practical performance of PatchAD on two real-world datasets. PatchAD exhibits the capability to detect the majority of anomalies, albeit with two anomalies remaining undetected. Furthermore, a noticeable disparity is observed between the reconstructed features by PatchAD and the original features, indicating its proficiency in learning the underlying patterns within the data, rather than simply replicating the original features.
Figure 6: The case studies on SWAT and WADI datasets.
IV-C3 Scalability Analysis
[…]
V Mechanism and Theory
This section endeavors to analyze the mechanisms of PatchAD from both experimental and theoretical perspectives, further elucidating the principles underlying its effectiveness. It aims to clarify the design motivations behind PatchAD.
V-A Unveiling the Mechanism
Figure 7 illustrates the knowledge acquired by PatchAD in four distinct cases. Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of [1,2,5][1,2,5][ 1 , 2 , 5 ]. The Original represents the original data, while Rec 1 and Rec 2 correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively. It is observed that the Intra Mixer captures the generalized knowledge in the data, which encompasses the overall trend information. Conversely, the Inter Mixer learnsthe detailed features of the data, compensating for any aspects that may not be captured by the Intra Mixer. A comparison between case 1 and case 2 reveals that using three distinct patch scales facilitates the acquisition of more detailed representations and enhances the generalization of the acquired features. Note that the y-axis ranges for Figure 7(a) and Figure 7(b) are different. Moreover, these two mixers collaborate to segregate and amplify the distinctive features effectively. The decoupling of features assists PatchAD in amplifying the anomaly-related characteristics, as indicated in Figure 3(c), thereby enhancing the detectability of anomalies. The superior performance of PatchAD can also be attributed to the varying sizes of the Inter and Intra Mixer branches, which prevent overfitting. This is the reason why PatchAD does not require the construction of negative samples in contrastive learning, as the two different networks construct implicit positive and negative samples.
Moreover, we empirically validate this conclusion through experiments. Case 1 was chosen as the baseline, and we investigated the entropy of the features acquired by the Inter and Intra Mixers under varying patch size conditions. Figures 8(a) and 8(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process. This observation suggests that the models effectively acquire knowledge and reduce uncertainty. Furthermore, we noticed that the Inter Mixer demonstrates the highest learning speed and exhibits the most substantial decrease in entropy when the patch size is configured as [1, 2, 5]. Additionally, as the patch size decreases, the entropy of the Inter Mixer increases, and it actually rises when the patch size is set to 1. This implies that the network is required to learn an excessive number of features. Figure 8(c) computes the ratio of Intra to Inter entropy, and our findings indicate that as the patch size increases, the features acquired by the Intra Mixer gain greater significance, thereby accelerating model learning through the utilization of multi-scale patch sizes.
Figure 9 illustrates the visualization of feature distributions in the SWAT dataset under three conditions. It is observed that anomalies are more easily detected in the Intra Mixer view (See Figure 9(b)), as its distinctiveness is more pronounced compared to the other two views (Figure 9 (a) & (c)).
According to our analysis, maximizing the disparities between the Inter Mixer and Intra Mixer can enhance the model’s detection capabilities. During network optimization, the Inter-Intra loss compels the Intra Mixer to acquire more generalized representations, effectively mitigating model overfitting.
Figure 7: Visualization of the features learnt by Inter and Intra Mixer.
Figure 8: Entropy analysis.
Figure 9: Latent distribution’s visualization by T-SNE.
V-B Theory Analysis
Considering spatial constraints and readability, further details are elaborated in Appendix -B.
VI Conclusion
We presents PatchAD, an innovative algorithm tailored for time series anomaly detection. PatchAD developed a lightweight multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning. This architecture is unique in using a purely MLP-based approach to discern inter-patch and intra-patch relationships in time series sequences. By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns. Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation. Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation. The results demonstrate that proposed lightweight PatchAD outperforms over 30 methods across eight benchmark datasets from different application scenarios. Additionally, we have conducted a comprehensive and extensive analysis of PatchAD from both theoretical and experimental perspectives, confirming the alignment between the model’s structure and its design motivations.